### v1

# Marco Randoa1, Martin Jamesb, Alessandro Verria, Lorenzo Rosascoa, Agnese Seminarab1

Abstract
We consider the problem of olfactory searches in a turbulent environment. We focus on agents that respond solely to odor stimuli, with no access to spatial perception nor prior information about the odor location. We ask whether navigation strategies to a target can be learned robustly within a sequential decision making framework. We develop a reinforcement learning algorithm using a small set of interpretable olfactory states and train it with realistic turbulent odor cues. By introducing a temporal memory, we demonstrate that two salient features of odor traces, discretized in few olfactory states, are sufficient to learn navigation in a realistic odor plume. Performance is dictated by the sparse nature of turbulent plumes. An optimal memory exists which ignores blanks within the plume and activates a recovery strategy outside the plume. We obtain the best performance by letting agents learn their recovery strategy and show that it is mostly casting cross wind, similar to behavior observed in flying insects. The optimal strategy is robust to substantial changes in the odor plumes, suggesting minor parameter tuning may be sufficient to adapt to different environments.
aMalGa - DIBRIS, University of Genova, ItalybMalGa - DICCA, University of Genova, ItalyThe authors declare no conflict of interest.1To whom correspondence should be addressed. E-mail: marco.rando@edu.unige.it, agnese.seminara@unige.itAuthor’s contributions. M.R., A.V., L.R. and A.S. designed research; M.R., M.J., A.V., L.R. and A.S. performed research; M.R., L.R. and A.S. wrote manuscript.Keywords: Navigation |||| Reinforcement learning |||| Olfaction |||| Turbulence |||| Memory
Bacterial cells localize the source of an attractive chemical even if they hold no spatial perception. They respond solely to temporal changes in chemical concentration and the result of their response is that they move toward attractive stimuli by climbing concentration gradients [1]. Larger organisms also routinely sense chemicals in their environment to localize or escape targets, but cannot follow chemical gradients since turbulence breaks odors into sparse pockets and gradients lose information [2, 3, 4, 5, 6]. The question of which features of turbulent odor traces are used by animals for navigation is natural, but not well understood. Beyond olfaction, some animals could use also prior spatial information to navigate  [7, 8, 9, 10], but if and how chemosensation and spatial perception are coupled is still not clear.
An algorithmic perspective to olfactory navigation in turbulence can shed light on some of these questions. Without aiming at an exhaustive taxonomy, see e.g. [11] for a recent review, we recall some approaches relevant to put our contribution in context. One class of methods are biomimetic algorithms, where explicit navigation rules are crafted taking inspiration from animal behavior. An advantage of these methods is interpretability, in the sense that they provide insights into features that effectively achieve turbulent navigation, for example: odor presence/absence [12, 13, 14, 5]; odor slope at onset of detection [15]; number of detections in a given interval of time [16] and the time of odor onset [17]. On the flip side, in biomimetic algorithms behaviors are hardwired and typically reactive, not relying on any optimality criterion.
A way to tackle this shortcoming is to cast olfactory navigation within a sequential decision making framework [18]. In this context, navigation is formalized as a task with a reward for success; by maximizing reward, optimal strategies can be sought to efficiently reach the target. A byproduct is that most algorithmic choices can often be done in a principled way. Within this framework, some approaches make explicit use of spatial information. Bayesian algorithms use a spatial map to guess the target location and use odor to refine this guess or “belief”. A prominent algorithm for olfactory navigation based on the concept of belief is the information-seeking algorithm [3] akin to exploration heuristics widely used in robotics [19, 20] (see e.g. [21, 22]). Using Bayesian sequential decision making and the notion of beliefs, navigation can be formalized as a Partially Observable Markov Decision Process (POMDP) [23, 24, 25], that can be approximatively solved [26, 27, 28]. POMDP approaches are appealing since beliefs are a sufficient statistics for the entire history of odor detections. However, they are computationally cumbersome. Further, they leave the question open of whether navigation as sequential decision making can be performed using solely olfactory information.
Recently, two algorithms studied navigation as a response to olfactory input alone [29, 30]. In [29] artificial neural networks were shown to learn near optimal strategies, but they were trained on odor cues with limited sparsity, and training with sparse odor cues typical of turbulence remains to be tested. In [30] an approach based on finite state controllers was proposed. Here, optimization was done using a model-based technique, relying on prior knowledge of the likelihood to detect the odor in space, hence still using spatial information. A different model free optimization could also be considered avoiding spatial information but this latter approach also remains to be tested. More generally, all the above approaches manipulate internally the previous history (memory) of odor detections. In this sense they are less interpretable, since the features of odor traces that drive navigation do not emerge explicitly.
Figure 1: Learning a stimulus-response strategy for turbulent navigation. (A) Representation of the search problem with turbulent odor cues obtained from Direct Numerical Simulations of fluid turbulence (grey scale, odor snapshot from the simulations). The discrete position $s$ is hidden; the odor concentration $z_{T}=z(s(t^{\prime}),t^{\prime})|t-T\leq t^{\prime}\leq t$ is observed along the trajectory $s(t^{\prime})$, where $T$ is the sensing memory. (B) Odor traces from direct numerical simulations at different (fixed) points within the plume. Odor is noisy and sparse, information about the source is hidden in the temporal dynamics. (C) Contour maps of olfactory states with nearly infinite memory ($T=2598$): on average olfactory states map to different locations within the plume and the void state is outside the plume. (D) Performance of stimulus-response strategies obtained during training, averaged over 500 episodes. We train using realistic turbulent data with memory $T=20$ and backtracking recovery.
In this paper, we propose a reinforcement learning (RL) approach to navigation in turbulence based on a set of interpretable olfactory features, with no spatial information, and highlight the role played by memory within this context. More precisely, we learn optimal strategies from data by training tabular Q learning [18] with realistic odor cues obtained from state-of-the-art Direct Numerical Simulations of turbulence. From the odor cues, we define features as moving averages of odor intensity and sparsity: the moving window is the temporal memory and naturally connects to the physics of turbulent odors. States are then obtained discretizing such features. Due to sparsity, agents may detect no odor within the moving window. We show there is an optimal memory minimizing the occurrence of this “void state”. The optimal memory scales with the blank time dictated by turbulence as it emerges from a trade off requiring that: (i) short blanks– typical of turbulent plumes– are ignored by responding to detections further in the past, and (ii) long blanks promptly trigger a recovery strategy to make contact with the plume again. We leverage these observations to tune the memory adaptively, by setting it equal to the previous blank experienced along an agent’s path. With this choice, the algorithm tests successfully in distinct environments, suggesting that tuning can be made robustly to enable generalization. The agent learns to surge upwind in most non-void states and to recover by casting crosswind in the absence of detections. Optimal agents limit encounters with the void state to a narrow band right at the edge of the plume. This suggests that the temporal odor features we considered effectively predict when the agent is exiting the plume and point to an intimate connection between temporal predictions and spatial navigation.
Figure 2: The optimal memory $T^{*}$. (A) Four measures of performance as a function of memory with backtracking recovery (solid line) show that the optimal memory $T^{*}=20$ maximizes average performance and minimizes standard deviation, except for the normalized time. Top: Averages computed over $10$ realizations of test trajectories starting from $43000$ initial positions (dash: results with adaptive memory). Bottom: standard deviation of the mean performance metrics for each initial condition (see Materials and Methods). (B) Average number of times agents encounter the void state along their path, $\langle N_{\emptyset}\rangle$, as a function of memory (top); cumulative average reward $\langle G\rangle$ is inversely correlated to $\langle N_{\emptyset}\rangle$ (bottom), hence the optimal memory minimizes encounters with the void. (C) Colormaps: Probability that agents at different spatial locations are in the void state at any point in time, starting the search form anywhere in the plume and representative trajectory of a successful searcher (green solid line) with memory $T=1$, $T=20$, $T=50$ (left to right). At the optimal memory agents in the void state are concentrated near the edge of the plume. Agents with shorter memories encounter voids throughout the plume; agents with longer memories encounter more voids outside of the plume as they delay recovery. In all panels, shades are $\pm$ standard deviation.
Significance
Finding mates or food in the presence of turbulence is challenging because odors constantly switch on and off unpredictably. As a result, it is unclear whether animals couple odor to other sources of information, what are the relevant features of odor stimuli and how they change according to the environment. A long history of bioinspired algorithms address this problem by crafting rules for navigation that mimic animal behavior: but can effective navigation be learned from the environment rather than set a priori? To address this question we train a reinforcement learning algorithm with realistic turbulent stimuli. Searchers learn to navigate by trial and error and respond solely to odor, with no further input. All computations are defined explicitly, enhancing interpretability. The upshot is that the algorithm identifies odor features as averages over a temporal scale (memory) dictated by the time between odor detections and thus by physics. There is no need to know physics beforehand, as memory can be adjusted based on experience. This approach naturally complements previous algorithms that use prior information and a map of space to plan navigation, rather than learn it from the environment. To what extent different animals plan vs learn to navigate remains to be understood.
Results
Background
[…]
Using time vs space to address partial observability.
In our problem however, the agent does not know where the source is, hence its position $s$ relative to the source, is unknown or “partially observed”. Instead, it can sense odor released by the target. In the language of RL, odor is an “observation” – but does it hold information about the position $s$? The answer is yes: several properties of odor stimuli depend on the distance from the source. However in the presence of turbulence, information lies in the statistics of the odor stimulus. Indeed, when odor is carried by a turbulent flow, it develops into a dramatically stochastic plume, i.e. a complex and convoluted region of space where the fluid is rich in odor molecules. Turbulent plumes break into structures that distort and expand while they travel away from their source and become more and more diluted [31, 4, 32, 6], see Figure 1A. As a consequence, an agent within the plume experiences intermittent odor traces that endlessly switch on (whiff) and off (blank) Figure 1B. The intensity of odor whiffs and how they are interleaved with blanks depends on distance from release, as dictated by physics [32]. Thus the upshot of turbulent transport is that the statistical properties of odor traces depend intricately on the position of the agent relative to the source. In other words, information about the state $s$ is hidden within the observed odor traces.
This positional information can be leveraged with a Bayesian approach that relies on guessing $s$, i.e. defining the probability distribution of the position, also called belief. This is the approach that has been more commonly adopted in the literature until now [26, 27, 28]. Note that because of the complexity of these algorithms, only relatively simple measures of the odor are computationally feasible, e.g. instantaneous presence/absence. Here we take a different model-free and map-free approach. Instead of guessing the current state $s$, we ignore the spatial position and respond directly to the temporal traces of the odor cues. Two other algorithms have been proposed to solve partial observability by responding solely to odor traces with recurrent neural networks [29] and finite state controllers [30] that manipulate implicitly the odor traces. Here instead we manipulate odor traces explicitly, by defining memory as a moving window and by crafting a small number of features of odor traces.
Features of odor cues: definition of discrete olfactory states and sensing memory.
To learn a response to odor traces, we first define a finite set of olfactory states, $o\in O$, so that they bear information about the location $s$. Defining the olfactory states is a challenge due to the dramatic fluctuations and irregularity of turbulent odor traces. To construct a fully interpretable low dimensional state space, we aim at a small number of olfactory states that bear robust information about $s$, i.e. for all values of $s$. We previously found that pairing features of sparsity as well as intensity of turbulent odor traces predicts robustly the location of the source for all $s$ [33]. Guided by these results, we use these two features extracted from the temporal history of odor detections to define a small set of olfactory states.
We proceed to define a function that takes as input the history of odor detections along an agent’s path and returns its current olfactory state. We indicate with $s(t)$ the (unknown) path of an agent, and with $z(s(t),t)$ the observations i.e. odor detections along its path. First, we define a sensing memory $T$ and we consider a short excerpt of the history of odor detections $z_{T}$ of duration $T$ prior to the current time $t$. Formally, $z_{T}(t):=\{z(s(t^{\prime}),t^{\prime})\,|\,t-T\leq t^{\prime}\leq t\}$. Second, we measure the average intensity $c$ (moving average of odor intensity over the time window $T$, conditioned to times when odor is above threshold), and intermittency $i$ (the fraction of time the odor is above threshold during the sensing window $T$). Both features $c$ and $i$ are described by continuous, positive real numbers. Third, we define 15 olfactory states by discretizing $i$ and $c$ in 3 and 5 bins respectively. Intermittency $i$ is bounded between 00 and $1$ and we discretize it in 3 bins by defining two thresholds (33% and 66%). The average concentration, $c$, is bounded between 0 and the odor concentration at the source, hence prior information on the source is needed to discretize $c$ using set thresholds. To avoid relying on prior information, we define thresholds of intensity as percentiles, based on a histogram that is populated online, along each agent’s path (see Materials and Methods). The special case where no odor is detected over $T$ deserves attention, hence we include it as an additional state named “void state” and indicate it with $o\equiv\emptyset$. When $T$ is sufficiently long, the resulting olfactory states map to different spatial locations (Figure 1C, with $T$ equal to the simulation time). Hence this definition of olfactory states can potentially mitigate the problem of partial observability using temporal traces, rather than spatial maps. But will these olfactory states with finite memory $T$ guide agents to the source?
Q learning: a map-less and model-free navigation to odor sources.
To answer this question, we trained tabular episodic Q learning [18]. In a nutshell, we use a simulator to place an agent at a random location in space at the beginning of each episode. The agent is not aware of its location in space, but it senses odor provided by the fluid dynamics simulator and thus can compute its olfactory state $o$, based on odor detected along its path in the previous $T$ sensing window. It then makes a move according to a set policy of actions $a\sim\pi_{0}(a|o)$. After the move, the simulator displaces the agent to its new location and relays the agent a negative reward $R=-\eta$ if it is not at the source and a positive reward $R=1$ if it reaches the source. The goal of RL is to find a policy of actions that maximizes the expected cumulative future reward $G=E_{\pi}(\sum_{t=0}^{\infty}\gamma^{t}R_{t+1})$ where the expectation is over the ensemble of trajectories and rewards generated by the policy from any initial condition. Because reward is only positive at the source, the optimal policy is the one that reaches the source as fast as possible. To further encourage the agent to reach the source quickly, we introduce a discount factor $\gamma<1$.
Episodes where the agent does not reach the source are ended after $H_{\text{max}}=5000$ with no positive reward. As it tries actions and receives rewards, the agent learns how good the actions are. This is accomplished by estimating the quality matrix $Q(o,a)$, i.e. the maximum expected cumulative reward conditioned to being in $o$ and choosing action $a$ at the present time: $Q(o,a)=\max_{\pi}E_{\pi}(\sum_{t=0}^{\infty}\gamma^{t}R_{t+1}|o_{t}=o,a_{t}=a)$. At each step, the agent improves its policy by choosing more frequently putatively good actions. Once the agent has a good approximation of the quality matrix, the optimal policy corresponds to the simple readout: $\pi^{*}(a|o)=\delta(a-a^{*}(o))$ where $a^{*}(o)=\arg\max_{a}Q(o,a)$, for non-void states $o\neq\emptyset$.
Recovery strategy. To fully describe the behavior of our Q-learning agents, we have to prescribe their policy from the void state $o\equiv\emptyset$. This is problematic because turbulent plumes are full of holes thus the void state can occur anywhere both within and outside the plume, Figure 1A. As a consequence, the optimal action $a^{*}(\emptyset)$ from the void state is ill defined. We address this issue by using a separate policy called “recovery strategy”. Inspired by path integration as defined in biology [34, 35, 36], we propose the backtracking strategy consisting in retracing the last $T_{a}$ steps after the agent lost track of the odor. If at the end of backtracking the agent is still in the void state, it activates Brownian motion. Backtracking requires that we introduce memory of the past $T_{a}$ actions. This timescale $T_{a}$ for activating recovery is conceptually distinct from the duration of the sensing memory – however here we set $T_{a}=T$ for simplicity.
We find that Q-learning agents successfully learn to navigate to the odor source by responding solely to their olfactory state, with no sense of space nor models of the odor cues. Learning can be quantified by monitoring the cumulative reward which continuously improves with further training episodes (Figure 1D, left). Improved reward corresponds to agents learning how to reach the source more quickly and reliably with training. Indeed, it is easy to show that the expected cumulative reward $G=\langle e^{-\lambda\tau}-\eta(1-e^{-\lambda\tau})/(1-\gamma)\rangle$, where $\tau$ is a random variable corresponding to time to reach the source and $\gamma=e^{-\lambda\Delta t}$ is the discount factor, with the time step $\Delta t=1$ (see Materials and Methods). Large rewards arise when (i) a large fraction $f^{+}$ of agents successfully reaches the source and (ii) the agents reach the source quickly, which maximizes $g^{+}=\langle e^{-\lambda\tau}|\text{success}\rangle$. Indeed $G=f^{+}G^{+}+(1-f^{+})G^{-}$, where $G^{+}=g^{+}-\eta(1-g^{+})/(1-\gamma)$ and $G^{-}=-\eta(1-e^{-\lambda H_{\text{max}}})/(1-\gamma)$. $H_{\text{max}}$ is the horizon of the agent i.e. the maximum time the agent is allowed to search, and after which the search is considered failed. Note that agents starting closer to the target receive larger rewards purely because of their initial position. To monitor performance independently on the starting location, we introduce the inverse time to reach the source relative to the shortest-path time from the same initial location, which goes from 0 for failing agents to 1 for ideal agents $\langle\tau_{\text{min}}/\tau\rangle$, independently on their starting location. Note that this is not the quantity that is optimized for. One may specifically target this perfomance metrics, which is agnostic on the duration of an agent’s path, by discounting proportionally to $t/\tau_{\text{min}}$.
All four measures of performance plateau to a maximum, suggesting learning has achieved a nearly optimal policy (Figure 1D). Once training is completed, we simulate the trajectory of test-agents starting from any of the about 43 000 admissible locations within the plume and moving according to the optimal policy. We will recapitulate performance with the cumulative reward $G$ averaged over the test-agents and dissect it into speed $g^{+}$, convergence $f^{+}$ and relative time $\langle\tau_{\text{min}}/\tau\rangle$.
Figure 3: The adaptive memory approximates the duration of the blank dictated by physics and it is an efficient heuristics, especially when coupled with a learned recovery strategy. (A) Colormaps of the Eulerian blank times $\tau_{b}$ (top) and the sensing memory $T$ (bottom): Left: averages; Right: standard deviations. The sensing memory statistics is computed over all agents that are located at each discrete cell, at any point in time. (B) Probability distribution of $\tau_{b}$ across all spatial locations and times (black) and of $T$ across all agents at all times (gray). (C) Performance with the adaptive memory nears performance of the optimal fixed memory, here shown for backtracking; similar results apply to the Brownian recovery (Supplementary Figure 2). (D) Comparison of three recovery strategies with adaptive memory: The learned recovery with adaptive memory outperforms all fixed and adaptive memory agents. In (C) and (D) dark squares mark the mean, and light rectangles mark $\pm$ standard deviation. No standard deviation is shown for the $f^{+}$ measure in the learned case as this strategy is deterministic (see Materials and Methods).
Optimal memory.
By repeating training using different values of $T$ we find that performance depends on memory and an optimal memory $T^{*}$ exists (Figure 2A). Why is there an optimal memory? The shortest memory $T=1$ corresponds to instantaneous olfactory states: the instantaneous contour maps of the olfactory states are convoluted and the void state is pervasive (Figure 2C, top). As a consequence, agents often activate recovery even when they are within the plume, the policy almost always leads to the source ($f^{+}=79\%\pm 13\%$) but follows lengthy convoluted paths ($\tau_{\text{min}}/\tau=0.14\pm 0.05$, Figure 2A, bottom). As memory increases, the olfactory states become smoother and agents encounter less voids (Figure 2C, center), perform straighter trajectories ($\tau_{\text{min}}/\tau=0.5\pm 0.3$) and reach the source reliably ($f^{+}=95\%\pm 8\%$), Figure 2C, bottom. Further increasing memory leads to even less voids within the plume and even smoother olfactory states. However – perhaps surprisingly – performance does not further improve but slightly decreases (at $T=50$, $f^{+}=94\%\pm 8\%$ and $\tau_{\text{min}}/\tau=0.38\pm 0.36$). A long memory is deleterious because it delays recovery from accidentally exiting the plume, thus increases the number of voids outside of the plume (Figure 2C, bottom). Indeed, agents often leave the plume accidentally as they measure their olfactory state while they move. They receive no warning, but realize their mistake after $T$ steps, when they enter the void state and activate recovery to re-enter the plume. The delay is linear with memory when agents recover by backtracking, but it depends on the recovery strategy (see discussion below and Supplementary Figure 1).
Thus short memories increase time in void within the plume, whereas long memories increase time in void outside the plume: the optimal memory minimizes the overall chances to experience the void (Figure 2B). Intuitively, $T^{*}$ should match the typical duration $\tau_{b}$ of blanks encountered within the plume, so that voids within the plume are effectively ignored without delaying recovery unnecessarily. Consistently, $\langle\tau_{b}\rangle$ averaged across all locations and times within the plume is $\langle\tau_{b}\rangle=9.97\pm 41.16$, comparable with the optimal memory $T^{*}$ (Figure 2A).
Adaptive memory.
There is no way to select the optimal memory $T^{*}$ without comparing several agents or relying on prior information on the blank durations. In order to avoid prior information, we venture to define memory adaptively along each agent’s path, using the intuition outlined above. We define a buffer memory $T_{b}$, and let the agent respond to a sensing window $T<T_{b}$. Ideally we would like to set $T\sim\langle\tau_{b}\rangle$. With this choice, blanks shorter than the average blank are ignored, as they are expected within the plume, whereas blanks longer than average initiate recovery, as they signal that the agent exited the plume. However, agents do not have access to $\langle\tau_{b}\rangle$ hence we set $T=\tau_{b}^{-}$, where $\tau_{b}^{-}$ is the most recent blank experienced by the agent. With this choice, the sensing memory $T$ fluctuates considerably along an agent’s path, due to turbulence ([32] and Figure 3A-B). Note that blanks are estimated along paths, thus the statistics of $T$ only qualitatively matches the Eulerian statistics of $\tau_{b}$. Despite the fluctuations, performance using the adaptive memory nears performance with the optimal memory (Figure 3C). This result confirms our intuition that memory should match the blank time. The advantage of the adaptive memory is that it relies solely on experience, with no prior information whatsoever. This is unlike $T^{*}$ which can only be selected using prior information, with no guarantee of generalization to other plumes.
Figure 4: Optimal policies with adaptive memory for different recovery strategies: backtracking (green), Brownian (red) and learned (blue). For each recovery, we show the spatial distribution of the olfactory states (top); the policy (center) and the state occupancy (bottom) for non-void states (left) vs the void state $\pi^{*}(a|\emptyset)$(right). Spatial distribution: probability that an agent at a given position is in any non-void olfactory state (left) or in the void state (right), color-coded from yellow to blue. Policy: actions learned in the non-void states $\sum_{o\neq\emptyset}n_{o}\pi^{*}(a|o)$, weighted on their occupancy $n_{o}$ (left, arrows proportional to the frequency of the corresponding action) and schematic view of recovery policy in the void state (right). State occupancy: fraction of agents that is in any of the 15 non-void states (left) or in the void state (right) at any point in space and time. Occupancy is proportional to the radius of the corresponding circle. The position of the circle identifies the olfactory state (rows and columns indicate the discrete intensity and intermittency respectively). All statistics is computed over 43000 trajectories, starting from any location within the plume.
Learning to recover.
So far, our agents combine a learned policy from non-void states to a heuristics from the void state, which we called the recovery strategy. We have considered a biologically-inspired heuristics where searchers make it back to locations within the plume by retracing their path backward. Similar results are obtained for Brownian recovery with a different optimal memory (see Materials and Methods and Supplementary Figure 1). To further strip the algorithm of heuristics, we ask whether the recovery strategy may be learned, rather than fixed a priori. To this end, we split the void state in many states, labeled with the time elapsed since first entering the void. The counter is reset to 00 whenever the searcher detects the odor. The definition of the 15 non-void states $o_{1},...,o_{15}$ remains unaltered. Interestingly, with this added degree of freedom, the agent learns an even better recovery strategy as reflected by all our measures of performance Figure 3D. Note that the learned recovery strategy resembles the casting behavior observed in flying insects [37], as discussed below.
Figure 5: Generalization to statistically different environments. (A) Snapshots of odor concentration normalizes with concentration at the source, colorcoded from blue (0) to yellow (1) for environment 1 to 6 from top to bottom. Environment $1^{*}$ is the native environment where all agents are trained. (B) Performance for the three recovery strategies backtracking (green) Brownian (red) and learned (blue), with adaptive memory, trained on the native environment and tested across all environments. Four measures of performance defined in the main text are shown. Dark squares mark the mean, and empty rectangles $\pm$ standard deviation. No standard deviation is shown for the $f^{+}$ measure in the learned case as this strategy is deterministic (see Materials and Methods).
Characterization of the optimal policies. To understand how different recoveries affect the agent’s behavior, we characterize the optimal policies obtained using the three recovery strategies. We visualize the probability to encounter each of the 16 olfactory states, or occupancy (circles in Figure 4), and the spatial distribution of the olfactory states. In the void state, the agent activates the recovery strategy. Recovery from the void state affects non-void olfactory states as well: their occupancy, their spatial distribution, and the action they elicit (Figure 4 and Supplementary Figure 3). This is because the agent computes its olfactory state online, according to its prior history which is affected by encounters with the void state. However, for all recoveries, non-void states are mostly encountered within the plume and largely elicit upwind motion (Figure 4, top, center). Thus macroscopically, all agents learn to surge upwind when they detect any odor within their memory, and to recover when their memory is empty. This suggests a considerable level of redundancy which may be leveraged to reduce the number of olfactory states, thus the computational cost.
The void state shows the most relevant differences: for both heuristic recoveries, 40% or more of the agents are in the void state and they are spatially spread out. In contrast, in the case of learned recovery, the optimal policy limits occurrence of the void state to 26% of the agents, confined to a narrow band near the edge of the plume. From these locations, the agents quickly recover the plume, explaining the boost in performance discussed above. Note that, exclusively for the learned recovery, the optimal policy is enriched in actions downwind to avoid overshooting the source. Indeed, from positions beyond the source, the learned strategy is unable to recover the plume as it mostly casts sideways, with little to no downwind action. Intuitively, the precise locations where agents move downwind may be crucial to efficiently avoid overshooting. Thus the policy may depend on specific details of the odor plume, consistent with poorer generalization of the learned recovery (discussed next).
Tuning for adaptation to different environments.
Finally, we test performance of the trained agents on six environments, characterized by distinct fluid flows and odor plumes (Figure 5 and Materials and Methods). Environment 1 is the native environment, where the agents were originally trained; Environment 2 is obtained by increasing the threshold of detection, which makes the signals considerably more sparse with longer blanks. Environments 3 and 4 are closer to the lower surface of the simulated domain, where the plume is smaller and fluctuates less. Environment 5 is a similar geometry, but obtained for a smaller Reynolds number and a different way to generate turbulence. Finally Environment 6 has an even larger Reynolds number, a longer domain and a smaller source, which creates an even more dramatically sparse signal. We consider agents with adaptive memory and compare the three recovery strategies discussed above – Brownian, backtracking and learned, see Figure 5B. Comparing performance across environments we find that: (i) although performance is degraded when testing in non native environments, all agents with adaptive memory are still extremely likely to find the source; (ii) Brownian recovery has lowest performance and generalization across all environments; (iii) backtracking provides good performance and generalization; (iii) the learned recovery strategy performs best in all environments by all performance metrics. In the most intermittent Environment 6 a striking 91% of agents succeeds in finding the source, with trajectories less the twice as long as the shortest path to the source. The upshot of generalization is that agents may navigate distinct turbulent plumes using a baseline strategy learned in a specific plume. Importantly, even if performance (mildly) degrades, most agents still do reach the source, suggesting that fine-tuning this strategy may enable efficient adaptation to different environments. Further work is needed to establish how much fine-tuning is needed to fully adapt the baseline strategy to different environments.
Discussion
In this work, we showed that agents exposed to a turbulent plume learn to associate salient features of the odor time trace – the olfactory state – to an optimal move that guides them to the odor source. The upshot of responding solely to odor is that the agent does not navigate based on where it believes the target is and thus needs no map of space nor prior information about the odor plume, which avoids considerable computational burden. On the flip side, in our stimulus-response algorithm, agents need to start from within the plume, however sparse and fragmented. Indeed, far enough from the source, Q-learning agents are mostly in the void state and they can only recover the plume if they have previously detected the odor or are right outside the plume. In contrast, agents using a map of space can navigate from larger distances than are reachable by responding directly to odor cues. Indeed, in the map-based POMDP setting, absence of odor detection is still informative and it enables agents to first find the plume, and then refine the search to localize the target within the plume [26, 38].
We show that because turbulent odor plumes constantly switch on and off, navigation must handle both absence and presence of odor stimuli. We address this fundamental issue by alternating between two distinct strategies: (i) Prolonged absence of odor prompts entry in the void state and triggers a recovery strategy to make contact with the plume again. We explored two heuristic recoveries and found that back-tracking to where the agent last detected odor is much more efficient than Brownian recovery. An even more efficient recovery can be learned that resembles cross-wind casting and limits the void state to a narrow region right outside of the plume. Casting is a well-studied computational strategy [12, 5] also observed in animal behavior, most famously in flying insects [37]. Intriguingly, cast and surge also emerges in algorithms making use of a model of the odor, whether for Bayesian updates or for policy optimization [3, 26, 30]. Whether natural casting behavior is learned, as in Q-learning, or is hard-wired in a model of the odor plume remains a fascinating question for further research. (ii) Odor detections prompt entry in non-void olfactory states, which predominantly elicit upwind surge. Blanks shorter than the sensing memory are ignored, i.e. agents do not enact recovery but respond to stimuli experienced prior to the short blank. Further work may optimize these non-void olfactory states by feature engineering, e.g. testing different discretizations to reduce redundancy or screening a large library of features using supervised learning as in [33] to potentially improve performance. Alternatively, feature engineering may be bypassed altogether by the use of recurrent neural networks (RNNs) as recently proposed in [29], possibly at the expense of interpretability. A systematic comparison using a common dataset is needed to elucidate how other heuristic and normative model-free algorithms handle odor presence vs odor absence.
To switch between the odor-driven strategy and the recovery strategy, we introduce a timescale $T$, which is an explicit form of temporal memory. $T$ delimits a sensing window extending in the recent past, prior to the present time. All odor stimuli experienced within the sensing window affect the current response. By using fixed memories of different duration, we demonstrate that an optimal memory exists and that the optimal memory minimizes the occurrence of the void state. On the one hand, long memories are detrimental because they delay recovery from accidentally exiting the plume. On the other hand, short memories are detrimental because they trigger recovery unnecessarily, i.e. even for blanks typically experienced within the turbulent plume. The optimal memory thus matches the typical duration of the blanks. To avoid using prior information on the statistics of the odor, we propose a simple heuristics setting memory adaptively equal to the most recent blank experienced along the path. The adaptive memory nears optimal performance despite dramatic fluctuations dictated by turbulence. Success of the heuristics suggests that a more accurate estimate of the future blank time may enable an even better adaptive memory; further work is needed to corroborate this idea.
Thus in Q-learning, memory is a temporal window matching odor blanks and distinguishing whether agents are in or out of the plume. The role of memory for olfactory search has been recently discussed in ref. [30]. In POMDPs, memory is stored in a detailed belief of agent position relative to the source. In finite state controllers, memory denotes an internal state of the agent and was linked to a coarse grained belief of the searcher being within or outside of the plume, similar to our findings. In recurrent neural networks memory is stored in the learned weights. A quantitative relationship between these different forms of memory and their connection to spatial perception remains to be understood.
We conclude by listing a series of experiments to test these ideas in living systems. First, olfactory search in living systems displays memory ([30, 10] and refs. therein). In insects, temporal scales can be measured associated to memory. Indeed, for flying insects loss of contact with a pheromone plume triggers crosswind casting and sometimes even downwind displacement [7, 39]. Interestingly, the onset of casting is delayed with respect to loss of contact with the plume, but this delay is not understood [39, 40]. In walking flies, the timing of previous odor encounters biases navigation [41]. (How) do these temporal timescales depend on the waiting times between previous detections? Using optogenetics [42, 43, 44, 45] or olfactory virtual reality with controlled odor delivery [46] experiments may measure memory as a function of the full history of odor traces. For insects, one may monitor memory by tracking the onset of cross wind casting with respect to the loss of the plume. More in general, a temporal memory may be defined by monitoring how far back in the past should two odor traces be identical in order to elicit the same repertoire of motor controls.
Second, our algorithm learns a stimulus-response strategy that relies solely on odor cues. The price to pay is that the agent must follow the ups and downs of the odor trace in order to compute averages and recognize blanks. A systematic study may use our algorithm to test the requirements of fidelity of this temporal representation, and how it depends on turbulence. How does turbulence affect the fidelity of odor temporal representation in living systems? Crustaceans provide excellent model system to ask this question, as they are known to use bursting olfactory receptor neurons to encode temporal information from olfactory scenes [47, 48]. Temporal information is also encoded in the olfactory bulb of mammals [49, 50]. Organisms with chemo-tactile systems like the octopus [51] may serve as a comparative model, to ask whether touch-chemosensation displays a sloppier temporal response, reflecting that surface-bound stimuli are not intermittent.
Third, our Q-learning algorithm requires the agent to receive olfactory information, thus start near or within the odor plume. In contrast, algorithms making use of a spatial map and prior information on the odor plume may first search for the plume (in conditions of near zero information) and then search the target within the plume [26, 38, 3]. Animals are known to use prior information to home into regions of space where the target is more likely to be found; but they can switch to navigation in response to odor (see e.g. [7, 8, 9, 10]). What triggers the switch from navigation driven by spatial perception to navigation driven by odor? For mice, the need for spatial perception may be tested indirectly by comparing paths in light vs dark, noting that neuronal place fields, that mediate spatial perception, are better stabilized by vision than olfaction [52, 53]. Thus in the light, animals have the ability to implement both map-less and map-based algorithms, whereas in the dark they are expected to more heavily rely on map-less algorithms. To make sure animals start searching for the odor target even before sensing odor, operant conditioning can be deployed so that animals associate an external cue (e.g. a sound) to the beginning of the task.
The reinforcement learning view of olfactory navigation offers an exciting opportunity to probe how living systems interact with the environment to accomplish complex real-world tasks affected by uncertainty. Coupling time varying odor stimuli with spatial perception is an instance of the broader question asking how animals combine prior knowledge regarding the environment with reaction to sensory stimuli. We hope that our work will spark further progress into connecting these broader questions to the physics of fluids.
Acknowledgements
This research was supported by grants to A.S. from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101002724 RIDING), the Air Force Office of Scientific Research under award number FA8655-20-1-7028, and the National Institutes of Health (NIH) under award number R01DC018789. L.R. acknowledges the financial support of the European Research Council (grant SLING 819789). We thank Francesco Viola for support and discussions regarding computational fluid dynamics as well as Antonio Celani, Venkatesh Murthy, Yujia Qi, Francesco Boccardo, Luca Gagliardi, Francesco Marcolli and Arnaud Ruymaekers for comments on the manuscript.
Materials and Methods
Data description.
The data we used to train the agents is a set of $2598$ matrices $\{D_{t}\}_{t=1}^{2598}$. Every matrix $D_{t}\in\mathbb{R}^{1225\times 280}$ contains the odor intensity in every position $(i,j)$ i.e. $(D_{t})_{i,j}$ represents the odor intensity in position $(i,j)$ at time $t$. The source of odor is in position $(20,142)$ and, in order to simplify the training, we considered as terminal states every position in a circle centered in the source position and with radius $10$ called source region. Data are obtained from a direct numerical simulation of the Navier-Stokes equations and the equations of transport of the odor. Environments 1 to 4 are derived from simulations of a channel flow described in ref [33]; Environment 5 corresponds to an additional simulation described below. We preprocess the data to eliminate simulation noise by setting to zero every entry of these matrices when they are smaller than a noise level $n_{\text{lvl}}:=3\times 10^{-6}$. Data information are summarized in Table 1.
Table 1: Data information
Feature Value
Shape $1225\times 280$
Number of time slices $2598$
$n_{\text{thr}}$ $3\times 10^{-6}$
Source location $(20,142)$
In environment 5, the odor is advected by a turbulent open channel flow, with three hemispherical obstacles placed on the ground close to the inlet to generate turbulence. The Navier-Stokes equations (1) and advection-diffusion equation for odor transport (2) are solved using a central second order finite difference scheme. The convective terms are discretized in time using an explicit Adams – Bashforth method; and the viscous and diffusion terms using an implicit Crank-Nicolson method [54]. The code is written in Fortran and is GPU parallelized. The channel is divided into $1024\times 256\times 128$ grid points along streamwise, spanwise and wall-normal directions respectively. The corresponding average spatial resolutions are $\Delta x=5\eta,\Delta y=5\eta,\Delta z=4\eta$, where $\eta$ is the Kolmogorov length scale. Three hemispheres of radius $100\eta$ are placed at a distance of $250\eta$ from the inlet on the ground, equally spaced along the spanwise direction. The channel is forced using a constant pressure gradient. For the velocity field, we impose a no-slip boundary condition at the ground and on the obstacles (${\bm{u}}=0$) and a free-slip boundary on top ($u_{z}=0,\partial_{z}u_{x}=\partial_{z}u_{y}=0$). The velocity field is periodic along the streamwise and spanwise directions. The bulk Reynolds number is 7800. For the odor field, we impose Dirichlet condition ($c=0$) at the ground, on the obstacles and inlet, no-flux ($\partial_{z}c=0$) on top, and outflow along other directions. Similar to the native environment, we choose the Schmidt number to be 1. The odor source is located downstream of the obstacle and centered at [$640\eta$, $640\eta$, $256\eta$] along streamwise, spanwise and wall-normal directions respectively. The odor source has a Gaussian profile with a standard deviation of $8\eta$.
Environment 6 is similar to environment 5 albeit with a higher bulk Reynolds number of 17500. Here, the channel is divided into $2000\times 500\times 200$ grid points and has an average spatial resolution of $\Delta x=\Delta y=\Delta z=5.5\eta$. The odor source has a Gaussian profile centered at [$825\eta$, $1375\eta$, $550\eta$] with a standard deviation of $3\eta$.
$\rho(\frac{\partial\bm{u}}{\partial t}+\bm{u}\cdot\nabla\bm{u})$ $=-\nabla P+\mu\nabla^{2}\bm{u}+\bm{f};$ (1)
$\nabla\cdot\bm{u}$ $=0.$
$\frac{\partial c}{\partial t}+\mathbf{u}\cdot\nabla c=D\nabla^{2}c+s.$ (2)
Olfactory states, Features & Discretization.
Each agent stores the odor concentrations detected in the previous $T$ time steps in a vector $\mathbf{M}$. We introduce an adaptive sensitivity threshold function $s_{\text{thr}}(\cdot)$ defined as
$s_{\text{thr}}(T):=\max\Big{\{}\frac{C_{\text{thr}}}{T}\sum\limits_{i=1}^{T}M_{i},n_{\text{thr}}\Big{\}},$ (3)
where $M_{i}$ denotes the $i$-th element of $M$ and $C_{\text{thr}}>0$ is a scaling constant (in our experiments we set it as $0.5$). $T$ denotes the cardinality of $M$. Given a memory $M$, we can define the filtered memory $\Delta^{M}$ as the set which contains every element of the memory $M$ that is higher than the sensitivity threshold $s_{\text{thr}}(M)$ i.e.
$\Delta^{M}:=\{z\in M\,|\,z>s_{\text{thr}}(M)\}.$ (4)
Then at timestep $t$, given the agent memory $M_{t}$, we define the average intensity $c(M_{t})$ and the intermittency $i(M_{t})$ as:
$c(M_{t})$ :=$:=\left\{\begin{array}[]{ll}\frac{1}{|\Delta^{M_{t}}|}\sum\limits_{i=1}^{|\Delta^{M_{t}}|}\Big{(}\Delta^{M_{t}}\Big{)}_{i},&|\Delta^{M_{t}}|>0\\ &\\ &\\ 0&\\ \end{array}\right.,$ (5)
$i(M_{t})$ $:=\frac{|\Delta^{M_{t}}|}{|M_{t}|}.$
Note that the average intensity is defined on the filtered memory $\Delta^{M}$, i.e. conditioned to detecting odors above threshold. Since the features defined in (5) returns real numbers, in order to use (tabular) q-learning, we need to discretize them. We denote with $\bar{i}(M_{t})$ the discretized intermittency. This is defined as follow
$\bar{i}(M_{t})$ :=$:=\left\{\begin{array}[]{ll}0,&\text{if}\quad i(M_{t})\leq 0.33\\ 1,&\text{if}\quad 0.33<i(M_{t})\leq 0.66\\ 2,&\text{if}\quad i(M_{t})>0.66\end{array}\right..$ (6)
The average intensity is bounded between zero and the maximum concentration of odor at the source. To avoid prior information on the source, we use a more structured procedure to discretize the average intensity online, based on the agent’s experience only. At every timestep $t$, the average intensity $c(M_{t})$ is computed and collected in a dataset $X_{t}$ i.e.
$X_{t}:=\{c(M_{0}),\cdots,c(M_{t})\}.$
Then, its discretized value is obtained by the following rule:
$\bar{c}(M_{t},X_{t})$ :=$:=\left\{\begin{array}[]{ll}0,&c(M_{t})\leq\text{p}(X_{t},25)\\ 1,&\text{p}(X_{t},25)<c(M_{t})\leq\text{p}(X_{t},50)\\ 2,&\text{p}(X_{t},50)<c(M_{t})\leq\text{p}(X_{t},80)\\ 3,&\text{p}(X_{t},80)<c(M_{t})\leq\text{p}(X_{t},99)\\ 4,&c(M_{t})>\text{p}(X_{t},99)\end{array}\right.,$ (7)
where $p(X_{t},n)$ denotes the $n$-th percentile of $X_{t}$. Finally, we can define the feature map $\phi_{t}$ as a function of the memory $M_{t}$ and the dataset of average intensities $X_{t}$ at timestep $t$
$\phi_{t}(M_{t},X_{t}):=[\bar{i}(M_{t}),\bar{c}(M_{t},X_{t})].$
This defines the current olfactory state $s_{t}$ i.e. at timestep $t$, the agent is in the olfactory state $o_{t}:=\phi_{t}(M_{t},X_{t})$. The case where the agent has no odor detections above threshold in its current memory, i.e. $|\Delta(M_{t})|=0$ corresponds to an additional state called void state ($\emptyset$) in the main text.
Agent Behavior and Policies.
Now, we describe how the agent interacts with the environment to solve the navigation problem. At every time step $t\in\mathbb{N}$, the agent observes an odor point $z_{t}$ and updates its memory including the new observation and removing the oldest i.e. it defines a memory $M_{t}$ with the following rule
$M_{t}:=\Bigg{[}\Big{(}M_{t-1}\Big{)}_{2},\cdots,\Big{(}M_{t-1}\Big{)}_{|M_{t-1}|},o_{t}\Bigg{]}.$ (8)
Then, it updates the dataset of average intensities i.e. $X_{t}:=X_{t-1}\cup\{c(M_{t})\}$ and it computes the olfactory state $o_{t}$. According to $o_{t}$, the agent chooses an action $a_{t}$ using a policy. As indicated in the main text, actions are the coordinate directions i.e. we define an action set $\mathcal{A}$ as follow
$\mathcal{A}:=\{e_{1},e_{2},-e_{1},-e_{2}\},$
where $e_{i}$ denotes the $i$-th canonical base. As explained in the main text, actions are selected using one of two policies according to the current olfactory state $o_{t}$. More precisely, if the olfactory state $o_{t}$ is not the void state, then the ($\epsilon$-greedy) Q-learning policy is used. Formally, let $Q$ be the Q matrix of the agent and let $o_{t}\neq\emptyset$, then the agent plays the action $a_{t}$ such that
$a_{t}=\left\{\begin{array}[]{ll}a\in\arg\max\limits_{a\in\mathcal{A}}Q(o_{t},a)&\text{with probability }1-\epsilon\\ a\sim\mathcal{U}(\mathcal{A})&\text{with probability }\epsilon\end{array}\right.,$ (9)
where, with $a\sim\mathcal{U}(\mathcal{A})$, we indicate an action $a$ uniformly sampled from $\mathcal{A}$. At test phase, the exploration-exploitation parameter $\epsilon$ is set to 00 and, thus, in an olfactory state $o_{t}\neq\emptyset$ the policy is deterministic. While training phase behavior is described in next paragraphs. In the void state $o_{t}=\emptyset$, the agent chooses the action $a_{t}\in\mathcal{A}$ according to a separated policy called recovery strategy. In our experiments, we defined and compared three different recovery strategies: Brownian, Backtracking and Learned.
Brownian recovery.
It is the simplest strategy we consider, consisting of playing random actions in the void state. Suppose that at time step $t$, the agent is in the void olfactory state, i.e., $o_{t}=\emptyset$, then $a_{t}$ is sampled uniformly from the action set $\mathcal{A}$. However, it is important to note for long memories agents start to recover when they are already far from the plume, and hitting the plume by random walk is prohibitively long. To avoid wandering away from the plume, the memory is constrained to be shorter, consistent with the observation that the optimal memory is $T^{*}=3$ to $5$, much shorter than for backtracking. At this memory, several blanks within the plume will cause the agent to recover, hence the lower performance of the Brownian recovery.
Backtracking Recovery.
In order to accelerate recovery from accidentally exiting the plume, we let the agents backtrack to the position where they last detected the odor. To this end, we first enumerate the actions with numbers from one to four. Then we introduce a new memory called action memory $A$. For simplicity, we consider the setting in which $|A|=|M|$. At time-step $t=0$, this memory is initialized as a vector of zeros indicating that the action memory is empty i.e. we define $A_{0}\in\mathbb{N}^{|M|}$ such that for every $i=1,\cdots,|A|$
$A_{i}=0.$
For every timestep $t>0$, the agent observes an odor point $z_{t}$ and updates the memory through (8). Moreover, the action memory is updated according to the status of the memory. If the last observation is smaller than the sensitivity threshold i.e. $z_{t}<s_{\text{thr}}(M_{t})$, the action previously played $a_{t-1}$ (represented by a natural number in $[1,4]$) is stored in the action memory i.e. for some $\Delta>0$, let
$A_{t-1}=[a_{t-\Delta},\cdots,a_{t-2},0,\cdots,0].$
Then
$A_{t}=[a_{t-\Delta},\cdots,a_{t-2},a_{t-1},\cdots,0].$
If at time-step $t$, the observation $z_{t}$ is larger than the sensitivity threshold then the action memory is reset i.e. $A_{t}\in\mathbb{N}^{|M|}$ with $(A_{t})_{i}=0$ for every $i$. If at timestep $t$, the memory is empty i.e. $c(M_{t})=0$, then the backtracking procedure is executed: the last non-zero element of the action memory is extracted and the inverse action is played i.e. For some $\Delta>0$, let
$A_{t-1}=[a_{t-\Delta},\cdots,a_{t-2}].$
Then, it plays the action $a_{t-2}$ and updates the action memory as follow
$A_{t}=[a_{t-\Delta},\cdots,a_{t-3},0].$
This procedure is repeated until either an observation larger than the sensitivity threshold is obtained or the action memory becomes empty. In the former case, the action memory is cleared and the action is chosen according to the Q-learning policy ((9)). In the latter case, a random action is played.
Note that this strategy only provides exploration after the backtracking fails to recover detections. Also, if agents start with no detection at time 0, the procedure is equivalent to Brownian motion.
Learned recovery.
In this case recovery policy is learned by splitting the void state in several states labeled by the time since entry in the void state. In our experiments, we split the void state in 30 states. Actions are then learned as in all other non-void states and the optimal action is always chosen with (9).
Training
An agent start at a random location within the odor plume at time 00. Its memory is initialized with the prior $|M_{0}|$ odor detections at its initial location $M_{0}=[z_{-|M_{0}|},\cdots,z_{0}]$, obtained from the fluid dynamics simulation. The Q-function $Q_{0}$ is initialized with $0.6$ for all actions and olfactory states. The first dataset of average intensities contains the first value $X_{0}=\{c(M_{0})\}$. At every times step $t>0$, the agent gets an odor observation $z_{t}$ from its new position and updates its memory including the new observation and removing the oldest and the olfactory state $o_{t}$ is computed (as described in previous paragraphs). The dataset of average intensities is updated: $X_{t}=X_{t-1}\cup\{c(M_{t})\}$. Exploration-exploitation parameter $\epsilon_{k}$ is scheduled as follow
$\epsilon_{k}=\eta_{\text{init}}\exp(-\eta_{\text{decay}}k),$
where, in our experiments, $\eta_{\text{init}}=0.99$ and $\eta_{\text{decay}}=0.0001$. At every episode $k$, the Q-function is updated at every time step $t$ as
$Q_{k+1}(s_{t},a_{t}):=(1-\alpha_{k})Q_{k}(s_{t},a_{t})+\alpha_{k}(r_{t}+\gamma\max\limits_{a^{\prime}}Q_{k}(s_{t+1},a^{\prime})),$
where $R_{t}$ is the immediate reward received playing the action $a_{t}$. $o_{t}$ and $o_{t+1}$ are the current and the next olfactory states and $\alpha_{k}$ is the learning rate at episode $k$. This is scheduled as
$\alpha_{k}=\alpha_{\text{init}}\exp(-\alpha_{\text{decay}}k),$
where, in our experiments, $\alpha_{\text{init}}=0.25$ and $\alpha_{\text{decay}}=0.001$. For the experiments, agents are trained in $100000$ episodes and an horizon of $5000$ steps. The agent velocity is set to $10$ and the discount factor is $\gamma=0.9999$.
Agents Evaluation.
To evaluate the performance of the different agents, we consider four metrics: the cumulative reward $G$ (which is the actual quantity that the algorithm optimizes for); normalized time (defined below); the fraction of success $f^{+}$ and the value conditioned on success $g^{+}$. For a fixed position $(i,j)$, we denote with $\tau_{\text{min}}(i,j)$ the minimum number of steps required to reach the source region from $(i,j)$ i.e. the length of the shortest path.
We define $D_{\text{init}}$ the set of points in which the first observation is above the sensitivity threshold (valid points). For each initial position $(i,j)\in D_{\text{init}}$, let $\tau(i,j)$ be the duration of the path obtained by an agent to reach the source. Note that $\tau(i,j)$ is a random variable for the stochastic backtracking and Brownian recoveries, but it is deterministic for the learned strategy that has no random components. For each admissible location $(i,j)$, we define four performance metrics:
$G(i,j)=$ $\langle e^{-\lambda\tau(i,j)}-\frac{\eta}{1-\gamma}(1-e^{-\lambda\tau(i,j)})\rangle$
$f^{+}(i,j)=$ $\frac{n_{\text{success}}(i,j)}{n_{\text{reps}}}$
$g^{+}(i,j)=$ $\langle e^{-\lambda\tau(i,j)}|\text{success}\rangle$
$\frac{\tau_{\text{min}}}{\tau}(i,j)=$ $\langle\frac{\tau_{\text{min}}(i,j)}{\tau(i,j)}\rangle$
[…]

### v2

@@
-Marco Randoa1, Martin Jamesb, Alessandro Verria, Lorenzo Rosascoa, Agnese Seminarab1
+Q-learning with temporal memory to navigate turbulence
 Abstract
 We consider the problem of olfactory searches in a turbulent environment.
-We focus on agents that respond solely to odor stimuli, with no access to spatial perception nor prior information about the odor location.
-We ask whether navigation strategies to a target can be learned robustly within a sequential decision making framework.
+We focus on agents that respond solely to odor stimuli, with no access to spatial perception nor prior information about the odor.
+We ask whether navigation to a target can be learned robustly within a sequential decision making framework.
 We develop a reinforcement learning algorithm using a small set of interpretable olfactory states and train it with realistic turbulent odor cues.
 By introducing a temporal memory, we demonstrate that two salient features of odor traces, discretized in few olfactory states, are sufficient to learn navigation in a realistic odor plume.
-Performance is dictated by the sparse nature of turbulent plumes.
+Performance is dictated by the sparse nature of turbulent odors.
 An optimal memory exists which ignores blanks within the plume and activates a recovery strategy outside the plume.
@@
 The optimal strategy is robust to substantial changes in the odor plumes, suggesting minor parameter tuning may be sufficient to adapt to different environments.
-aMalGa - DIBRIS, University of Genova, ItalybMalGa - DICCA, University of Genova, ItalyThe authors declare no conflict of interest.1To whom correspondence should be addressed.
-E-mail: marco.rando@edu.unige.it, agnese.seminara@unige.itAuthor’s contributions.
-M.R., A.V., L.R.
-and A.S.
-designed research; M.R., M.J., A.V., L.R.
-and A.S.
-performed research; M.R., L.R.
-and A.S.
-wrote manuscript.Keywords: Navigation |||| Reinforcement learning |||| Olfaction |||| Turbulence |||| Memory
+1 Introduction
 Bacterial cells localize the source of an attractive chemical even if they hold no spatial perception.
-They respond solely to temporal changes in chemical concentration and the result of their response is that they move toward attractive stimuli by climbing concentration gradients [1].
-Larger organisms also routinely sense chemicals in their environment to localize or escape targets, but cannot follow chemical gradients since turbulence breaks odors into sparse pockets and gradients lose information [2, 3, 4, 5, 6].
+They respond solely to temporal changes in chemical concentration and the result of their response is that they move toward attractive stimuli by climbing concentration gradients Berg (1975).
+Larger organisms also routinely sense chemicals in their environment to localize or escape targets, but cannot follow chemical gradients since turbulence breaks odors into sparse pockets and gradients lose information Murlis et al.
+(1992); Vergassola et al.
+(2007); Shraiman and Siggia (2000); Balkovsky and I.
+(2002); Reddy et al.
+(2022).
 The question of which features of turbulent odor traces are used by animals for navigation is natural, but not well understood.
-Beyond olfaction, some animals could use also prior spatial information to navigate  [7, 8, 9, 10], but if and how chemosensation and spatial perception are coupled is still not clear.
+Beyond olfaction, some animals could use also prior spatial information to navigate  Cardé (2021); Schal (1982); Gire et al.
+(2016); Baker et al.
+(2018), but if and how chemosensation and spatial perception are coupled is still not clear.
 An algorithmic perspective to olfactory navigation in turbulence can shed light on some of these questions.
 Without aiming at an exhaustive taxonomy, see e.g.
-[11] for a recent review, we recall some approaches relevant to put our contribution in context.
+Celani and Panizon (2024) for a recent review, we recall some approaches relevant to put our contribution in context.
 One class of methods are biomimetic algorithms, where explicit navigation rules are crafted taking inspiration from animal behavior.
-An advantage of these methods is interpretability, in the sense that they provide insights into features that effectively achieve turbulent navigation, for example: odor presence/absence [12, 13, 14, 5]; odor slope at onset of detection [15]; number of detections in a given interval of time [16] and the time of odor onset [17].
+An advantage of these methods is interpretability, in the sense that they provide insights into features that effectively achieve turbulent navigation, for example: odor presence/absence Baker (1990); Kramer (1997); Belanger and Willis (1988); Balkovsky and I.
+(2002); odor slope at onset of detection Atema (1996); number of detections in a given interval of time Michaelis et al.
+(2020) and the time of odor onset Demir et al.
+(2020a).
 On the flip side, in biomimetic algorithms behaviors are hardwired and typically reactive, not relying on any optimality criterion.
-A way to tackle this shortcoming is to cast olfactory navigation within a sequential decision making framework [18].
+A way to tackle this shortcoming is to cast olfactory navigation within a sequential decision making framework Sutton and Barto (1998).
 In this context, navigation is formalized as a task with a reward for success; by maximizing reward, optimal strategies can be sought to efficiently reach the target.
@@
 Bayesian algorithms use a spatial map to guess the target location and use odor to refine this guess or “belief”.
-A prominent algorithm for olfactory navigation based on the concept of belief is the information-seeking algorithm [3] akin to exploration heuristics widely used in robotics [19, 20] (see e.g.
-[21, 22]).
-Using Bayesian sequential decision making and the notion of beliefs, navigation can be formalized as a Partially Observable Markov Decision Process (POMDP) [23, 24, 25], that can be approximatively solved [26, 27, 28].
+A prominent algorithm for olfactory navigation based on the concept of belief is the information-seeking algorithm Vergassola et al.
+(2007) akin to exploration heuristics widely used in robotics Cassandra et al.
+(1996); LaValle (2006) (see e.g.
+Loisy and Eloy (2022); Ishida (2020)).
+Using Bayesian sequential decision making and the notion of beliefs, navigation can be formalized as a Partially Observable Markov Decision Process (POMDP) Krishnamurthy (2016); Hauskrecht (2000); Shani et al.
+(2013), that colorblackcan be approximatively solved Rigolli et al.
+(2022b); Heinonen et al.
+(2023); Loisy and Heinonen (2023b).
 POMDP approaches are appealing since beliefs are a sufficient statistics for the entire history of odor detections.
@@
 Further, they leave the question open of whether navigation as sequential decision making can be performed using solely olfactory information.
-Recently, two algorithms studied navigation as a response to olfactory input alone [29, 30].
-In [29] artificial neural networks were shown to learn near optimal strategies, but they were trained on odor cues with limited sparsity, and training with sparse odor cues typical of turbulence remains to be tested.
-In [30] an approach based on finite state controllers was proposed.
-Here, optimization was done using a model-based technique, relying on prior knowledge of the likelihood to detect the odor in space, hence still using spatial information.
+Recently, two algorithms studied navigation as a response to olfactory input alone Singh et al.
+(2023); Verano et al.
+(2023).
+In Singh et al.
+(2023) artificial neural networks were shown to learn near optimal strategies as a response to odor and instantaneous flow direction, although they were trained on odor cues with limited sparsity, and training with sparse odor cues typical of turbulence remains to be tested.
+In Verano et al.
+(2023) an approach based on finite state controllers was proposed.
+Here, optimization was done assuming fixed known mean flow direction, and using a model-based technique, relying on prior knowledge of the likelihood to detect the odor in space, hence still using spatial information.
 A different model free optimization could also be considered avoiding spatial information but this latter approach also remains to be tested.
@@
 In this sense they are less interpretable, since the features of odor traces that drive navigation do not emerge explicitly.
-Figure 1: Learning a stimulus-response strategy for turbulent navigation.
-(A) Representation of the search problem with turbulent odor cues obtained from Direct Numerical Simulations of fluid turbulence (grey scale, odor snapshot from the simulations).
-The discrete position $s$ is hidden; the odor concentration $z_{T}=z(s(t^{\prime}),t^{\prime})|t-T\leq t^{\prime}\leq t$ is observed along the trajectory $s(t^{\prime})$, where $T$ is the sensing memory.
-(B) Odor traces from direct numerical simulations at different (fixed) points within the plume.
-Odor is noisy and sparse, information about the source is hidden in the temporal dynamics.
-(C) Contour maps of olfactory states with nearly infinite memory ($T=2598$): on average olfactory states map to different locations within the plume and the void state is outside the plume.
-(D) Performance of stimulus-response strategies obtained during training, averaged over 500 episodes.
-We train using realistic turbulent data with memory $T=20$ and backtracking recovery.
-In this paper, we propose a reinforcement learning (RL) approach to navigation in turbulence based on a set of interpretable olfactory features, with no spatial information, and highlight the role played by memory within this context.
-More precisely, we learn optimal strategies from data by training tabular Q learning [18] with realistic odor cues obtained from state-of-the-art Direct Numerical Simulations of turbulence.
+In this paper, we propose a reinforcement learning (RL) approach to navigation in turbulence based on a set of interpretable olfactory features, with no spatial information other than the ability to orient relative to the mean flow, and highlight the role played by memory within this context.
+More precisely, we learn optimal strategies from data by training tabular Q learning Sutton and Barto (1998) with realistic odor cues obtained from state-of-the-art Direct Numerical Simulations of turbulence.
 From the odor cues, we define features as moving averages of odor intensity and sparsity: the moving window is the temporal memory and naturally connects to the physics of turbulent odors.
@@
 This suggests that the temporal odor features we considered effectively predict when the agent is exiting the plume and point to an intimate connection between temporal predictions and spatial navigation.
-Figure 2: The optimal memory $T^{*}$.
-(A) Four measures of performance as a function of memory with backtracking recovery (solid line) show that the optimal memory $T^{*}=20$ maximizes average performance and minimizes standard deviation, except for the normalized time.
-Top: Averages computed over $10$ realizations of test trajectories starting from $43000$ initial positions (dash: results with adaptive memory).
-Bottom: standard deviation of the mean performance metrics for each initial condition (see Materials and Methods).
-(B) Average number of times agents encounter the void state along their path, $\langle N_{\emptyset}\rangle$, as a function of memory (top); cumulative average reward $\langle G\rangle$ is inversely correlated to $\langle N_{\emptyset}\rangle$ (bottom), hence the optimal memory minimizes encounters with the void.
-(C) Colormaps: Probability that agents at different spatial locations are in the void state at any point in time, starting the search form anywhere in the plume and representative trajectory of a successful searcher (green solid line) with memory $T=1$, $T=20$, $T=50$ (left to right).
-At the optimal memory agents in the void state are concentrated near the edge of the plume.
-Agents with shorter memories encounter voids throughout the plume; agents with longer memories encounter more voids outside of the plume as they delay recovery.
-In all panels, shades are $\pm$ standard deviation.
-Significance
-Finding mates or food in the presence of turbulence is challenging because odors constantly switch on and off unpredictably.
-As a result, it is unclear whether animals couple odor to other sources of information, what are the relevant features of odor stimuli and how they change according to the environment.
-A long history of bioinspired algorithms address this problem by crafting rules for navigation that mimic animal behavior: but can effective navigation be learned from the environment rather than set a priori?
-To address this question we train a reinforcement learning algorithm with realistic turbulent stimuli.
-Searchers learn to navigate by trial and error and respond solely to odor, with no further input.
-All computations are defined explicitly, enhancing interpretability.
-The upshot is that the algorithm identifies odor features as averages over a temporal scale (memory) dictated by the time between odor detections and thus by physics.
-There is no need to know physics beforehand, as memory can be adjusted based on experience.
-This approach naturally complements previous algorithms that use prior information and a map of space to plan navigation, rather than learn it from the environment.
-To what extent different animals plan vs learn to navigate remains to be understood.
-Results
+Figure 1: Learning a stimulus-response strategy for turbulent navigation.
+(A) Representation of the search problem with turbulent odor cues obtained from Direct Numerical Simulations of fluid turbulence (grey scale, odor snapshot from the simulations).
+The discrete position $s$ is hidden; the odor concentration $z_{T}=z(s(t^{\prime}),t^{\prime})|t-T\leq t^{\prime}\leq t$ is observed along the trajectory $s(t^{\prime})$, where $T$ is the sensing memory.
+(B) Odor traces from direct numerical simulations at different (fixed) points within the plume.
+Odor is noisy and sparse, information about the source is hidden in the temporal dynamics.
+(C) Contour maps of olfactory states with nearly infinite memory ($T=2598$): on average olfactory states map to different locations within the plume and the void state is outside the plume.
+Intermittency is discretized in three bins defined by two thresholds 66% (red line) and 33% (blue line).
+Intensity is discretized in 5 bins (dark red shade to white shade) defined by four thresholds (percentiles 99%, 80%, 50%, 25%).
+(D) Performance of stimulus-response strategies obtained during training, averaged over 500 episodes.
+We train using realistic turbulent data with memory $T=20$ and backtracking recovery.
+2 Results
 Background
@@
 In the language of RL, odor is an “observation” – but does it hold information about the position $s$?
-The answer is yes: several properties of odor stimuli depend on the distance from the source.
+The answer is yes: several properties of odor stimuli depend on the distance from the source Boie et al.
+(2018); Ackels et al.
+(2021b); Nag and van Breugel (2024).
 However in the presence of turbulence, information lies in the statistics of the odor stimulus.
@@
 a complex and convoluted region of space where the fluid is rich in odor molecules.
-Turbulent plumes break into structures that distort and expand while they travel away from their source and become more and more diluted [31, 4, 32, 6], see Figure 1A.
+Turbulent plumes break into structures that distort and expand while they travel away from their source and become more and more diluted Falkovich et al.
+(2001); Shraiman and Siggia (2000); Celani et al.
+(2014); Reddy et al.
+(2022), see Figure 1A.
 As a consequence, an agent within the plume experiences intermittent odor traces that endlessly switch on (whiff) and off (blank) Figure 1B.
-The intensity of odor whiffs and how they are interleaved with blanks depends on distance from release, as dictated by physics [32].
+The intensity of odor whiffs and how they are interleaved with blanks depends on distance from release, as dictated by physics Celani et al.
+(2014).
 Thus the upshot of turbulent transport is that the statistical properties of odor traces depend intricately on the position of the agent relative to the source.
@@
 defining the probability distribution of the position, also called belief.
-This is the approach that has been more commonly adopted in the literature until now [26, 27, 28].
+This is the approach that has been more commonly adopted in the literature until now Rigolli et al.
+(2022b); Heinonen et al.
+(2023); Loisy and Heinonen (2023b).
 Note that because of the complexity of these algorithms, only relatively simple measures of the odor are computationally feasible, e.g.
@@
 Instead of guessing the current state $s$, we ignore the spatial position and respond directly to the temporal traces of the odor cues.
-Two other algorithms have been proposed to solve partial observability by responding solely to odor traces with recurrent neural networks [29] and finite state controllers [30] that manipulate implicitly the odor traces.
+Two other algorithms have been proposed to solve partial observability by responding solely to odor traces with recurrent neural networks Singh et al.
+(2023) and finite state controllers Verano et al.
+(2023) that manipulate implicitly the odor traces.
 Here instead we manipulate odor traces explicitly, by defining memory as a moving window and by crafting a small number of features of odor traces.
 Features of odor cues: definition of discrete olfactory states and sensing memory.
-To learn a response to odor traces, we first define a finite set of olfactory states, $o\in O$, so that they bear information about the location $s$.
+To learn a response to odor traces, we set out to craft a finite set of olfactory states, $o\in O$, so that they bear information about the location $s$.
 Defining the olfactory states is a challenge due to the dramatic fluctuations and irregularity of turbulent odor traces.
@@
 for all values of $s$.
-We previously found that pairing features of sparsity as well as intensity of turbulent odor traces predicts robustly the location of the source for all $s$ [33].
+We previously found that pairing features of sparsity as well as intensity of turbulent odor traces predicts robustly the location of the source for all $s$ Rigolli et al.
+(2022a).
 Guided by these results, we use these two features extracted from the temporal history of odor detections to define a small set of olfactory states.
@@
 We indicate with $s(t)$ the (unknown) path of an agent, and with $z(s(t),t)$ the observations i.e.
-odor detections along its path.
+odor concentration along its path.
 First, we define a sensing memory $T$ and we consider a short excerpt of the history of odor detections $z_{T}$ of duration $T$ prior to the current time $t$.
@@
 Q learning: a map-less and model-free navigation to odor sources.
-To answer this question, we trained tabular episodic Q learning [18].
+To answer this question, we trained tabular episodic Q learning Sutton and Barto (1998).
 In a nutshell, we use a simulator to place an agent at a random location in space at the beginning of each episode.
@@
 It then makes a move according to a set policy of actions $a\sim\pi_{0}(a|o)$.
-After the move, the simulator displaces the agent to its new location and relays the agent a negative reward $R=-\eta$ if it is not at the source and a positive reward $R=1$ if it reaches the source.
+After the move, the simulator displaces the agent to its new location and relays the agent a penalty $R=-\sigma$ with $\sigma=0.001$ if it is not at the source and a reward $R=1$ if it reaches the source.
 The goal of RL is to find a policy of actions that maximizes the expected cumulative future reward $G=E_{\pi}(\sum_{t=0}^{\infty}\gamma^{t}R_{t+1})$ where the expectation is over the ensemble of trajectories and rewards generated by the policy from any initial condition.
@@
 We address this issue by using a separate policy called “recovery strategy”.
-Inspired by path integration as defined in biology [34, 35, 36], we propose the backtracking strategy consisting in retracing the last $T_{a}$ steps after the agent lost track of the odor.
+Inspired by path integration as defined in biology Etienne and Jeffery (2004); Etienne et al.
+(1996); Heinze et al.
+(2018), we propose the backtracking strategy consisting in retracing the last $T_{a}$ steps after the agent lost track of the odor.
 If at the end of backtracking the agent is still in the void state, it activates Brownian motion.
@@
 This timescale $T_{a}$ for activating recovery is conceptually distinct from the duration of the sensing memory – however here we set $T_{a}=T$ for simplicity.
+Backtracking was observed in ants displaced in unfamiliar environments Wystrach et al.
+(2013), tsetse flies executing reverse turns bringing them back towards the location where they last detected odor Torr (1988); Gibson and Brady (1985) and cockroaches retracing their steps downwind, sometimes walking all the way back to the release point upon plume loss Willis et al.
+(2008); it was also previously used in computational models Park et al.
+(2016).
 We find that Q-learning agents successfully learn to navigate to the odor source by responding solely to their olfactory state, with no sense of space nor models of the odor cues.
@@
 Improved reward corresponds to agents learning how to reach the source more quickly and reliably with training.
-Indeed, it is easy to show that the expected cumulative reward $G=\langle e^{-\lambda\tau}-\eta(1-e^{-\lambda\tau})/(1-\gamma)\rangle$, where $\tau$ is a random variable corresponding to time to reach the source and $\gamma=e^{-\lambda\Delta t}$ is the discount factor, with the time step $\Delta t=1$ (see Materials and Methods).
+Indeed, it is easy to show that the expected cumulative reward $G=\langle e^{-\lambda\tau}-\sigma(1-e^{-\lambda\tau})/(1-\gamma)\rangle$, where $\tau$ is a random variable corresponding to time to reach the source and $\gamma=e^{-\lambda\Delta t}$ is the discount factor, with the time step $\Delta t=1$ (see Materials and Methods).
 Large rewards arise when (i) a large fraction $f^{+}$ of agents successfully reaches the source and (ii) the agents reach the source quickly, which maximizes $g^{+}=\langle e^{-\lambda\tau}|\text{success}\rangle$.
-Indeed $G=f^{+}G^{+}+(1-f^{+})G^{-}$, where $G^{+}=g^{+}-\eta(1-g^{+})/(1-\gamma)$ and $G^{-}=-\eta(1-e^{-\lambda H_{\text{max}}})/(1-\gamma)$.
+Indeed $G=f^{+}G^{+}+(1-f^{+})G^{-}$, where $G^{+}=g^{+}-\sigma(1-g^{+})/(1-\gamma)$ and $G^{-}=-\sigma(1-e^{-\lambda H_{\text{max}}})/(1-\gamma)$.
 $H_{\text{max}}$ is the horizon of the agent i.e.
@@
 Once training is completed, we simulate the trajectory of test-agents starting from any of the about 43 000 admissible locations within the plume and moving according to the optimal policy.
+Admissible locations are defined as any location where the odor is non-zero at least once within the entire simulation.
 We will recapitulate performance with the cumulative reward $G$ averaged over the test-agents and dissect it into speed $g^{+}$, convergence $f^{+}$ and relative time $\langle\tau_{\text{min}}/\tau\rangle$.
-Figure 3: The adaptive memory approximates the duration of the blank dictated by physics and it is an efficient heuristics, especially when coupled with a learned recovery strategy.
-(A) Colormaps of the Eulerian blank times $\tau_{b}$ (top) and the sensing memory $T$ (bottom): Left: averages; Right: standard deviations.
-The sensing memory statistics is computed over all agents that are located at each discrete cell, at any point in time.
-(B) Probability distribution of $\tau_{b}$ across all spatial locations and times (black) and of $T$ across all agents at all times (gray).
-(C) Performance with the adaptive memory nears performance of the optimal fixed memory, here shown for backtracking; similar results apply to the Brownian recovery (Supplementary Figure 2).
-(D) Comparison of three recovery strategies with adaptive memory: The learned recovery with adaptive memory outperforms all fixed and adaptive memory agents.
-In (C) and (D) dark squares mark the mean, and light rectangles mark $\pm$ standard deviation.
-No standard deviation is shown for the $f^{+}$ measure in the learned case as this strategy is deterministic (see Materials and Methods).
 Optimal memory.
+Figure 2: The optimal memory $T^{*}$.
+(A) Four measures of performance as a function of memory with backtracking recovery (solid line) show that the optimal memory $T^{*}=20$ maximizes average performance and minimizes standard deviation, except for the normalized time.
+Top: Averages computed over $10$ realizations of test trajectories starting from $43000$ initial positions (dash: results with adaptive memory).
+Bottom: standard deviation of the mean performance metrics for each initial condition (see Materials and Methods).
+(B) Average number of times agents encounter the void state along their path, $\langle N_{\emptyset}\rangle$, as a function of memory (top); cumulative average reward $\langle G\rangle$ is inversely correlated to $\langle N_{\emptyset}\rangle$ (bottom), hence the optimal memory minimizes encounters with the void.
+(C) Colormaps: Probability that agents at different spatial locations are in the void state at any point in time, starting the search form anywhere in the plume and representative trajectory of a successful searcher (green solid line) with memory $T=1$, $T=20$, $T=50$ (left to right).
+At the optimal memory agents in the void state are concentrated near the edge of the plume.
+Agents with shorter memories encounter voids throughout the plume; agents with longer memories encounter more voids outside of the plume as they delay recovery.
+In all panels, shades are $\pm$ standard deviation.
+[…]
+[Optimal memory with Brownian recovery]The role of temporal memory with Brownian recovery strategy (same as main Figure 2A).
+(A) Total cumulative reward (top left) and standard deviation (top right) as a function of memory showing an optimal memory $T^{*}=3$ for the Brownian agent.
+Other measures of performance with their standard deviations show the same optimal memory (bottom).
+The trade-off between long and short memories discussed in the main text holds, but here exiting the plume is much more detrimental because regaining position within the plume by Brownian motion is much lengthier.
+(B) As for the Backtracking agent, the optimal memory corresponds minimizes the number of times the agent encounters the void state.Refer to caption
 By repeating training using different values of $T$ we find that performance depends on memory and an optimal memory $T^{*}$ exists (Figure 2A).
@@
 The shortest memory $T=1$ corresponds to instantaneous olfactory states: the instantaneous contour maps of the olfactory states are convoluted and the void state is pervasive (Figure 2C, top).
-As a consequence, agents often activate recovery even when they are within the plume, the policy almost always leads to the source ($f^{+}=79\%\pm 13\%$) but follows lengthy convoluted paths ($\tau_{\text{min}}/\tau=0.14\pm 0.05$, Figure 2A, bottom).
-As memory increases, the olfactory states become smoother and agents encounter less voids (Figure 2C, center), perform straighter trajectories ($\tau_{\text{min}}/\tau=0.5\pm 0.3$) and reach the source reliably ($f^{+}=95\%\pm 8\%$), Figure 2C, bottom.
+As a consequence, agents often activate recovery even when they are within the plume, the policy almost always leads to the source ($f^{+}=79\%\pm 13\%$) but follows lengthy convoluted paths ($\tau_{\text{min}}/\tau=0.14\pm 0.05$, Figure 2C, bottom).
+As memory increases, the olfactory states become smoother and agents encounter less voids (Figure 2C, center), perform straighter trajectories ($\tau_{\text{min}}/\tau=0.5\pm 0.3$) and reach the source reliably ($f^{+}=95\%\pm 8\%$), Figure 2A, bottom.
 Further increasing memory leads to even less voids within the plume and even smoother olfactory states.
@@
 They receive no warning, but realize their mistake after $T$ steps, when they enter the void state and activate recovery to re-enter the plume.
-The delay is linear with memory when agents recover by backtracking, but it depends on the recovery strategy (see discussion below and Supplementary Figure 1).
+The delay is linear with memory when agents recover by backtracking, but it depends on the recovery strategy (see Materials and Methods and Figure 2–figure supplement 1).
 Thus short memories increase time in void within the plume, whereas long memories increase time in void outside the plume: the optimal memory minimizes the overall chances to experience the void (Figure 2B).
@@
 However, agents do not have access to $\langle\tau_{b}\rangle$ hence we set $T=\tau_{b}^{-}$, where $\tau_{b}^{-}$ is the most recent blank experienced by the agent.
-With this choice, the sensing memory $T$ fluctuates considerably along an agent’s path, due to turbulence ([32] and Figure 3A-B).
+With this choice, the sensing memory $T$ fluctuates considerably along an agent’s path, due to turbulence (Celani et al.
+(2014) and Figure 3A-B).
 Note that blanks are estimated along paths, thus the statistics of $T$ only qualitatively matches the Eulerian statistics of $\tau_{b}$.
@@
 This is unlike $T^{*}$ which can only be selected using prior information, with no guarantee of generalization to other plumes.
+Figure 3: The adaptive memory approximates the duration of the blank dictated by physics and it is an efficient heuristics, especially when coupled with a learned recovery strategy.
+(A) Top to bottom: Colormaps of the Eulerian average blank time $\tau_{b}$; average sensing memory $T$; standard deviation of Eulerian blank time and of sensing memory.
+The sensing memory statistics is computed over all agents that are located at each discrete cell, at any point in time.
+(B) Probability distribution of $\tau_{b}$ across all spatial locations and times (black) and of $T$ across all agents at all times (gray).
+(C) Performance with the adaptive memory nears performance of the optimal fixed memory, here shown for backtracking; similar results apply to the Brownian recovery (Figure 3–figure supplement 2).
+(D) Comparison of five recovery strategies with adaptive memory: The learned recovery with adaptive memory outperforms all fixed and adaptive memory agents.
+In (C) and (D) dark squares mark the mean, and light rectangles mark $\pm$ standard deviation.
+$f^{+}$ is defined as the fraction of agents that reach the target at test, hence has no standard deviation.
+[…]
+[All performance metrics with fixed and adaptive memory and two different recovery strategies.]All four measures of performance across agents with fixed memory and Backtracking vs Brownian recovery (green and red respectively, unframed boxes) and with adaptive memory for Backtracking, Brownian and Learned recovery (green, red and blue respectively, framed boxes)Refer to caption
+Learning to recover.
+So far, our agents combine a learned policy from non-void states to a heuristics from the void state, which we called the recovery strategy.
+We have considered a biologically-inspired heuristics where searchers make it back to locations within the plume by retracing their path backward.
+To further strip the algorithm of heuristics, we ask whether the recovery strategy may be learned, rather than fixed a priori.
+To this end, we split the void state in many states, labeled with the time elapsed since first entering the void.
+We pick 50 void states, as less than 50 void states results in no convergence, and states above 50 are useless because rarely visited.
+The counter is reset to 00 whenever the searcher detects the odor.
+The definition of the 15 non-void states $o_{1},...,o_{15}$ remains unaltered.
+Interestingly, with this added degree of freedom, the agent learns an even better recovery strategy as reflected by all our measures of performance Figure 3D.
+Note that the learned recovery strategy resembles the casting behavior observed in flying insects David et al.
+(1983), as discussed below.
+In fact, insects deploy a range of recovery strategies depending on locomotor mode and environment.
+To corroborate these results, we compare performance using two additional biologically-inspired recovery strategies, i.e.
+circling (observed in windless environments Stupski and van Breugel (2024)), and cast & surge David et al.
+(1983) as well as a Brownian recovery which does not have a direct biological relevance but represents a simple computational benchmark.
+The learned recovery outperforms all heuristic recoveries, as seen by the cumulative reward $G$ (Figure 3D).
+Circling is the second best recovery and shortly follows the learned recovery.
+Circling achieves nearly optimal performance by further decreasing failures (metrics $f^{+}$), but slowing down (metrics $g^{+}$ and $\tau_{min}/\tau$).
+Characterization of the optimal policies.
 Figure 4: Optimal policies with adaptive memory for different recovery strategies: backtracking (green), Brownian (red) and learned (blue).
@@
 All statistics is computed over 43000 trajectories, starting from any location within the plume.
-Learning to recover.
-So far, our agents combine a learned policy from non-void states to a heuristics from the void state, which we called the recovery strategy.
-We have considered a biologically-inspired heuristics where searchers make it back to locations within the plume by retracing their path backward.
-Similar results are obtained for Brownian recovery with a different optimal memory (see Materials and Methods and Supplementary Figure 1).
-To further strip the algorithm of heuristics, we ask whether the recovery strategy may be learned, rather than fixed a priori.
-To this end, we split the void state in many states, labeled with the time elapsed since first entering the void.
-The counter is reset to 00 whenever the searcher detects the odor.
-The definition of the 15 non-void states $o_{1},...,o_{15}$ remains unaltered.
-Interestingly, with this added degree of freedom, the agent learns an even better recovery strategy as reflected by all our measures of performance Figure 3D.
-Note that the learned recovery strategy resembles the casting behavior observed in flying insects [37], as discussed below.
-Figure 5: Generalization to statistically different environments.
-(A) Snapshots of odor concentration normalizes with concentration at the source, colorcoded from blue (0) to yellow (1) for environment 1 to 6 from top to bottom.
-Environment $1^{*}$ is the native environment where all agents are trained.
-(B) Performance for the three recovery strategies backtracking (green) Brownian (red) and learned (blue), with adaptive memory, trained on the native environment and tested across all environments.
-Four measures of performance defined in the main text are shown.
-Dark squares mark the mean, and empty rectangles $\pm$ standard deviation.
-No standard deviation is shown for the $f^{+}$ measure in the learned case as this strategy is deterministic (see Materials and Methods).
-Characterization of the optimal policies.
+[…]
+[…]
 To understand how different recoveries affect the agent’s behavior, we characterize the optimal policies obtained using the three recovery strategies.
@@
 In the void state, the agent activates the recovery strategy.
-Recovery from the void state affects non-void olfactory states as well: their occupancy, their spatial distribution, and the action they elicit (Figure 4 and Supplementary Figure 3).
+Recovery from the void state affects non-void olfactory states as well: their occupancy, their spatial distribution, and the action they elicit (Figure 4 and Figure 4–figure supplement 3).
 This is because the agent computes its olfactory state online, according to its prior history which is affected by encounters with the void state.
@@
 This suggests a considerable level of redundancy which may be leveraged to reduce the number of olfactory states, thus the computational cost.
-The void state shows the most relevant differences: for both heuristic recoveries, 40% or more of the agents are in the void state and they are spatially spread out.
-In contrast, in the case of learned recovery, the optimal policy limits occurrence of the void state to 26% of the agents, confined to a narrow band near the edge of the plume.
-From these locations, the agents quickly recover the plume, explaining the boost in performance discussed above.
+Reducing the number of non-empty olfactory states drastically to just 1 does indeed show degraded performance (see Figure 5 - figure supplement 5).
+A systematic optimization of odor representation requires a considerable reformulation of the algorithm which is beyond the scope of the current work.
 Note that, exclusively for the learned recovery, the optimal policy is enriched in actions downwind to avoid overshooting the source.
@@
 Thus the policy may depend on specific details of the odor plume, consistent with poorer generalization of the learned recovery (discussed next).
+We expect that in conditions where overshooting the source is more prominent, downwind motion may emerge as an effective component of the recovery strategy, similar to observations in insects (e.g.
+Wolf and Wehner (2000); Alvarez-Salvado et al.
+(2018).
+The void state shows the most relevant differences: for both heuristic recoveries, 40% or more of the agents are in the void state and they are spatially spread out.
+In contrast, in the case of learned recovery, the optimal policy limits occurrence of the void state to 26% of the agents, confined to a narrow band near the edge of the plume.
+From these locations, the agents quickly recover the plume, explaining the boost in performance discussed above.
+In all distinct trainings of the agents with learned-recovery, we observed that the trajectories in the void start with an initial surge of 6$\pm$2 steps; continue with either crosswind direction and then switch to the other side, with little to no downwind actions (see Figure 4 - Figure Supplement 4 and Table 1).
+This recovery thus mixes aspects of exploitation (surge) to aspects of exploration (cast): we defer a more in-depth analysis that disentangles these two aspects elsewhere.
+Table 1: Parameters of the learned recovery, statistics over 20 independent trainings.
+Initial surge upwind $6\pm 2$
+Total steps upwind $15\pm 2$
+Total steps downwind $1.3\pm 1.4$
+Total steps to the right $15\pm 3$
+Total steps to the left $18\pm 6$
 Tuning for adaptation to different environments.
@@
 Finally Environment 6 has an even larger Reynolds number, a longer domain and a smaller source, which creates an even more dramatically sparse signal.
-We consider agents with adaptive memory and compare the three recovery strategies discussed above – Brownian, backtracking and learned, see Figure 5B.
-Comparing performance across environments we find that: (i) although performance is degraded when testing in non native environments, all agents with adaptive memory are still extremely likely to find the source; (ii) Brownian recovery has lowest performance and generalization across all environments; (iii) backtracking provides good performance and generalization; (iii) the learned recovery strategy performs best in all environments by all performance metrics.
-In the most intermittent Environment 6 a striking 91% of agents succeeds in finding the source, with trajectories less the twice as long as the shortest path to the source.
+Environments 1, 2, 5 and 6 are representative of conditions encountered far from a substrate, e.g.
+by flying or swimming organisms.
+Environments 3 and 4 represent odor near surfaces relevant to terrestrial or benthic navigation (but not directly applicable to trail tracking, where odor traces on the substrate is tracked).
+Note that we consider a Schmidt number Sc = 1 appropriate for odors in air but not in water.
+However, we expect a weak dependence on the Schmidt number as the Batchelor and Kolmogorov scales are below the size of the source and we are interested in the large scale statistics Falkovich et al.
+(2001); Celani et al.
+(2014); Duplat et al.
+(2010).
+We consider agents with adaptive memory and compare the five recovery strategies discussed above – backtracking, learned, circling, cast & surge and Brownian, see Figure 5B.
+Comparing performance across environments we find that: (i) although performance is degraded when testing in non native environments, backtracking, learned and circling recoveries with adaptive memory are still extremely likely to find the source.
 The upshot of generalization is that agents may navigate distinct turbulent plumes using a baseline strategy learned in a specific plume.
-Importantly, even if performance (mildly) degrades, most agents still do reach the source, suggesting that fine-tuning this strategy may enable efficient adaptation to different environments.
-Further work is needed to establish how much fine-tuning is needed to fully adapt the baseline strategy to different environments.
-Discussion
+Importantly, as most of these agents still do reach the source, fine-tuning may enable efficient adaptation to different environments.
+Further work is needed to establish how much fine-tuning is needed to fully adapt to different environments.
+(ii) Brownian and cast& surge recoveries have lowest performance and generalization across all environments.
+Cast & surge is often used as a comparison Verano et al.
+(2023) and can be extremely effective: our results do not contradict the literature, but simply showcase that cast width and surge length need to be carefully defined.
+(iii) The cumulative reward $G$ shows that the learned recovery is the best at generalizing to environments 2, 5 and 6.
+Particularly, in the most intermittent Environment 6 a striking 91% of agents succeeds in finding the source, with trajectories less the twice as long as the shortest path to the source.
+(iv) Circling is the best at generalizing to environments 3 and 4, representative of less intermittent regions near the substrate.
+As observed for the native environment, circling favors success rate (metrics $f^{+}$) against speed (metrics $g^{+}$ and $\tau_{\text{min}}/\tau$).
+Figure 5: Generalization to statistically different environments.
+(A) Snapshots of odor concentration normalized with concentration at the source, colorcoded from blue (0) to yellow (1) for environment 1 to 6 as labeled.
+Environment $1^{*}$ is the native environment where all agents are trained.
+(B) Performance for the five recovery strategies brownian (red), backtracking (green), learned (blue), circling (orange) and zigzag (purple) with adaptive memory, trained on the native environment and tested across all environments 1 to 6.
+Four measures of performance defined in the main text are shown.
+Dark squares mark the mean, and empty rectangles $\pm$ standard deviation.
+No standard deviation is shown for the $f^{+}$ measure for the learned, circling and zigzag recoveries as these strategies are deterministic (see Materials and Methods).
+[…]
+3 Discussion
 In this work, we showed that agents exposed to a turbulent plume learn to associate salient features of the odor time trace – the olfactory state – to an optimal move that guides them to the odor source.
 The upshot of responding solely to odor is that the agent does not navigate based on where it believes the target is and thus needs no map of space nor prior information about the odor plume, which avoids considerable computational burden.
+The only spatial awareness needed to implement this algorithm is the ability to orient motion relative to the mean flow which is assumed known.
+In reality animals cannot measure the mean flow but rely on local measures of flow speed, using e.g.
+antennas for insects Reynolds et al.
+(2010), whiskers for rodents Yu et al.
+(2016) or the lateral line for marine organisms Liao (2006).
+Further work is needed to bridge the gap with our simplified setting.
 On the flip side, in our stimulus-response algorithm, agents need to start from within the plume, however sparse and fragmented.
@@
 In contrast, agents using a map of space can navigate from larger distances than are reachable by responding directly to odor cues.
-Indeed, in the map-based POMDP setting, absence of odor detection is still informative and it enables agents to first find the plume, and then refine the search to localize the target within the plume [26, 38].
-We show that because turbulent odor plumes constantly switch on and off, navigation must handle both absence and presence of odor stimuli.
+Indeed, in the map-based POMDP setting, absence of odor detection is still informative and it enables agents to first find the plume, and then refine the search to localize the target within the plume Rigolli et al.
+(2022b); Loisy and Heinonen (2023a).
+We show that because the odor signal within a turbulent plume constantly switches on and off, navigation must handle both absence and presence of odor stimuli.
 We address this fundamental issue by alternating between two distinct strategies: (i) Prolonged absence of odor prompts entry in the void state and triggers a recovery strategy to make contact with the plume again.
-We explored two heuristic recoveries and found that back-tracking to where the agent last detected odor is much more efficient than Brownian recovery.
+We explored four heuristic recoveries and found that spiralling around the location where the agent last detected odor is the most efficient heuristics, which privileges reliable success rather than speed.
 An even more efficient recovery can be learned that resembles cross-wind casting and limits the void state to a narrow region right outside of the plume.
-Casting is a well-studied computational strategy [12, 5] also observed in animal behavior, most famously in flying insects [37].
-Intriguingly, cast and surge also emerges in algorithms making use of a model of the odor, whether for Bayesian updates or for policy optimization [3, 26, 30].
+Casting is a well-studied computational strategy Baker (1990); Balkovsky and I.
+(2002) also observed in animal behavior, most famously in flying insects David et al.
+(1983).
+Intriguingly, cast and surge also emerges in algorithms making use of a model of the odor, whether for Bayesian updates or for policy optimization Vergassola et al.
+(2007); Rigolli et al.
+(2022b); Verano et al.
+(2023).
 Whether natural casting behavior is learned, as in Q-learning, or is hard-wired in a model of the odor plume remains a fascinating question for further research.
+Clearly, the width of the casts and length of the surges are of crucial importance: a hard-wired cast and surge recovery with arbitrary parameters shows poor performance.
 (ii) Odor detections prompt entry in non-void olfactory states, which predominantly elicit upwind surge.
@@
 agents do not enact recovery but respond to stimuli experienced prior to the short blank.
+The non-void olfactory states are crafted based on biologically-plausible features which have been shown to harness positional information.
 Further work may optimize these non-void olfactory states by feature engineering, e.g.
-testing different discretizations to reduce redundancy or screening a large library of features using supervised learning as in [33] to potentially improve performance.
-Alternatively, feature engineering may be bypassed altogether by the use of recurrent neural networks (RNNs) as recently proposed in [29], possibly at the expense of interpretability.
-A systematic comparison using a common dataset is needed to elucidate how other heuristic and normative model-free algorithms handle odor presence vs odor absence.
+testing different discretizations to reduce redundancy.
+A drastic reduction to a single non-void olfactory state degrades performance, suggesting screening a large library of features using supervised learning as in Rigolli et al.
+(2022a) may be used to potentially improve performance.
+This approach will provide a systematic ranking of the most efficient temporal features of odor time traces for navigation; however it will have to test different memories, discretizations and regression algorithms as well, making it cumbersome.
+Alternatively, feature engineering may be bypassed altogether by the use of recurrent neural networks (RNNs) Singh et al.
+(2023) or finite state controllers Verano et al.
+(2023).
+These algorithms are appealing in that they bypass entirely the need to hand-craft explicit features of odor time traces.
+On the flip side, they provide no explicit handle at the odor features that drive behavior nor at the specific duration of the temporal memory and how it is related to the physics of the odor cues.
+Thus to extract these information extra work is needed to interrogate these algorithms.
+For example, principal component analysis in Singh et al.
+(2023) suggests the hidden state of trained agents correlates with biologically relevant variables, including head direction, odor concentration and time since last detection.
+Finally, a systematic comparison using a common dataset is needed to elucidate how other heuristic and normative model-free algorithms handle odor presence vs odor absence.
 To switch between the odor-driven strategy and the recovery strategy, we introduce a timescale $T$, which is an explicit form of temporal memory.
@@
 The role of memory for olfactory search has been recently discussed in ref.
-[30].
+Verano et al.
+(2023).
 In POMDPs, memory is stored in a detailed belief of agent position relative to the source.
@@
 We conclude by listing a series of experiments to test these ideas in living systems.
-First, olfactory search in living systems displays memory ([30, 10] and refs.
+First, olfactory search in living systems displays memory (Verano et al.
+(2023); Baker et al.
+(2018) and refs.
 therein).
 In insects, temporal scales can be measured associated to memory.
-Indeed, for flying insects loss of contact with a pheromone plume triggers crosswind casting and sometimes even downwind displacement [7, 39].
-Interestingly, the onset of casting is delayed with respect to loss of contact with the plume, but this delay is not understood [39, 40].
-In walking flies, the timing of previous odor encounters biases navigation [41].
+Indeed, for flying insects loss of contact with a pheromone plume triggers crosswind casting and sometimes even downwind displacement Cardé (2021); Kuenen and Cardé (1994).
+Interestingly, the onset of casting is delayed with respect to loss of contact with the plume Kuenen and Cardé (1994); van Breugel F and MH (2014), but this delay is not understood.
+Similarly, an odor detection elicits upwind surges that can last several seconds Kathman et al.
+(2024); Alvarez-Salvado et al.
+(2018).
+In walking flies, the timing of previous odor encounters biases navigation Demir et al.
+(2020b).
 (How) do these temporal timescales depend on the waiting times between previous detections?
-Using optogenetics [42, 43, 44, 45] or olfactory virtual reality with controlled odor delivery [46] experiments may measure memory as a function of the full history of odor traces.
+Using optogenetics Gepner et al.
+(2015); Hernandez-Nunez et al.
+(2015); Matheson et al.
+(2022); David Stupski and van Breugel (2024) or olfactory virtual reality with controlled odor delivery Radvansky and Dombeck (2018) experiments may measure memory as a function of the full history of odor traces.
 For insects, one may monitor memory by tracking the onset of cross wind casting with respect to the loss of the plume.
@@
 How does turbulence affect the fidelity of odor temporal representation in living systems?
-Crustaceans provide excellent model system to ask this question, as they are known to use bursting olfactory receptor neurons to encode temporal information from olfactory scenes [47, 48].
-Temporal information is also encoded in the olfactory bulb of mammals [49, 50].
-Organisms with chemo-tactile systems like the octopus [51] may serve as a comparative model, to ask whether touch-chemosensation displays a sloppier temporal response, reflecting that surface-bound stimuli are not intermittent.
+Crustaceans provide excellent model system to ask this question, as they are known to use bursting olfactory receptor neurons to encode temporal information from olfactory scenes Bobkov and Ache (2007); Ache et al.
+(2016).
+Temporal information is also encoded in the olfactory bulb of mammals Carey et al.
+(2009); Ackels et al.
+(2021a).
+Organisms with chemo-tactile systems like the octopus Allard et al.
+(2023) may serve as a comparative model, to ask whether touch-chemosensation displays a sloppier temporal response, reflecting that surface-bound stimuli are not intermittent.
 Third, our Q-learning algorithm requires the agent to receive olfactory information, thus start near or within the odor plume.
-In contrast, algorithms making use of a spatial map and prior information on the odor plume may first search for the plume (in conditions of near zero information) and then search the target within the plume [26, 38, 3].
+In contrast, algorithms making use of a spatial map and prior information on the odor plume may first search for the plume (in conditions of near zero information) and then search the target within the plume Rigolli et al.
+(2022b); Loisy and Heinonen (2023a); Vergassola et al.
+(2007).
 Animals are known to use prior information to home into regions of space where the target is more likely to be found; but they can switch to navigation in response to odor (see e.g.
-[7, 8, 9, 10]).
+Cardé (2021); Schal (1982); Gire et al.
+(2016); Baker et al.
+(2018)).
 What triggers the switch from navigation driven by spatial perception to navigation driven by odor?
-For mice, the need for spatial perception may be tested indirectly by comparing paths in light vs dark, noting that neuronal place fields, that mediate spatial perception, are better stabilized by vision than olfaction [52, 53].
+For mice, the need for spatial perception may be tested indirectly by comparing paths in light vs dark, noting that neuronal place fields, that mediate spatial perception, are better stabilized by vision than olfaction Save et al.
+(2000); Zhang and D (2015).
 Thus in the light, animals have the ability to implement both map-less and map-based algorithms, whereas in the dark they are expected to more heavily rely on map-less algorithms.
@@
 a sound) to the beginning of the task.
+Note that distinct species control locomotion differently and as a result trajectories are usually far more complex than a sequence of discrete steps on a checker board.
+Thus to compare algorithms to animal behavior a more detailed model of the specific motor controls is to be developed.
 The reinforcement learning view of olfactory navigation offers an exciting opportunity to probe how living systems interact with the environment to accomplish complex real-world tasks affected by uncertainty.
@@
 We hope that our work will spark further progress into connecting these broader questions to the physics of fluids.
-Acknowledgements
-This research was supported by grants to A.S.
-from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101002724 RIDING), the Air Force Office of Scientific Research under award number FA8655-20-1-7028, and the National Institutes of Health (NIH) under award number R01DC018789.
-L.R.
-acknowledges the financial support of the European Research Council (grant SLING 819789).
-We thank Francesco Viola for support and discussions regarding computational fluid dynamics as well as Antonio Celani, Venkatesh Murthy, Yujia Qi, Francesco Boccardo, Luca Gagliardi, Francesco Marcolli and Arnaud Ruymaekers for comments on the manuscript.
-Materials and Methods
+4 Methods and Materials
 Data description.
-The data we used to train the agents is a set of $2598$ matrices $\{D_{t}\}_{t=1}^{2598}$.
+The data we used to train the agents (Environment 1) is a set of $2598$ matrices $\{D_{t}\}_{t=1}^{2598}$.
 Every matrix $D_{t}\in\mathbb{R}^{1225\times 280}$ contains the odor intensity in every position $(i,j)$ i.e.
@@
 Data are obtained from a direct numerical simulation of the Navier-Stokes equations and the equations of transport of the odor.
-Environments 1 to 4 are derived from simulations of a channel flow described in ref [33]; Environment 5 corresponds to an additional simulation described below.
-We preprocess the data to eliminate simulation noise by setting to zero every entry of these matrices when they are smaller than a noise level $n_{\text{lvl}}:=3\times 10^{-6}$.
-Data information are summarized in Table 1.
-Table 1: Data information
-Feature Value
-Shape $1225\times 280$
-Number of time slices $2598$
-$n_{\text{thr}}$ $3\times 10^{-6}$
-Source location $(20,142)$
-In environment 5, the odor is advected by a turbulent open channel flow, with three hemispherical obstacles placed on the ground close to the inlet to generate turbulence.
+Environments 1 to 4 are derived from Simulation 1, a direct numerical simulation of a channel flow described in Rigolli et al.
+(2022a) and used to develop a POMDP algorithm in Rigolli et al.
+(2022b), dataset available from Rigolli et al.
+(2022c).
+The simulation represents a boundary layer matching the lowest $\sim$1m of the atmosphere and with horizontal dimensions $\sim$1.9$\times$9.5 m and Reynolds number $Re_{\lambda}\sim 1400$ (on the low side of atmospheric Reynolds typically ranging from 1000 to 10000 Gulitski et al.
+(2007)).
+Odor snapshots are extracted at a height .$\sim.5\,m∼ .$ from the ground for Environments 1 and 2 and .$\sim.01\,m∼ .$ for Environments 3 and 4 respectively.
+We preprocess the data to zero every entry of these matrices when they are smaller than a noise level $n_{\text{lvl}}:=3\times 10^{-6}$ (or 0.13 % relative to concentration at the source).
+The noise level is increased to 0.22% in Environments 2 and 4.
+Data information are summarized in Table 2 and 3.
+Levels of intermittency in Figure 1 show that only a thin core region has intermittency larger than 66%, whereas the most challenging regions at the edge of the plume have intermittency under 33%.
+For reference, experimental values of 25% to 20% were reported for a surrogate odor in the atmospheric boundary layer, along the centerline at 2 to 15 m from the source Murlis and Jones (1981).
+Table 2: Gridworld geometry.
+From Top: 2D size of the simulation, agents that leave the simulation box continue to receive negative reward and no odor; number of time stamps in the simulation, beyond which simulations are looped; number of actions per time stamp; speed of the agent; noise level below which odor is not detected; location of the source on the grid.
+See Table 3 for the values of the grid size $\Delta x$ and time stamps at which odor snapshots are saved.
+Simulation 1 Simulation 2 Simulation 3
+2D simulation grid $1225\times 280$ $1024\times 256$ $2000\times 500$
+# time stamps $2598$ $5000$ $5000$
+# decisions per time stamp 1 1 1
+Speed (grid points / time stamp) 10 10 10
+$n_{\text{lvl}}$ $3\times 10^{-6}$ $3\times 10^{-6}$ $10^{-4}$
+Source location $(20,142)$ $(128,128)$ $(150,250)$
+Environments 5 and 6 correspond to horizontal slices at mid height extracted from two additional simulations we performed to corroborate the results (Simulation 2 and 3).
+In Simulation 2, the odor is advected by a turbulent open channel flow, with three hemispherical obstacles placed on the ground close to the inlet to generate turbulence.
 The Navier-Stokes equations (1) and advection-diffusion equation for odor transport (2) are solved using a central second order finite difference scheme.
-The convective terms are discretized in time using an explicit Adams – Bashforth method; and the viscous and diffusion terms using an implicit Crank-Nicolson method [54].
+The convective terms are discretized in time using an explicit Adams – Bashforth method; and the viscous and diffusion terms using an implicit Crank-Nicolson method Viola et al.
+(2020).
 The code is written in Fortran and is GPU parallelized.
@@
 The channel is forced using a constant pressure gradient.
-For the velocity field, we impose a no-slip boundary condition at the ground and on the obstacles (${\bm{u}}=0$) and a free-slip boundary on top ($u_{z}=0,\partial_{z}u_{x}=\partial_{z}u_{y}=0$).
+For the velocity field, we impose a no-slip boundary condition at the ground and on the obstacles (${\boldsymbol{u}}=0$) and a free-slip boundary on top ($u_{z}=0,\partial_{z}u_{x}=\partial_{z}u_{y}=0$).
 The velocity field is periodic along the streamwise and spanwise directions.
@@
 The odor source has a Gaussian profile with a standard deviation of $8\eta$.
-Environment 6 is similar to environment 5 albeit with a higher bulk Reynolds number of 17500.
+Simulation 3 is similar to Simulation 2, albeit with a higher bulk Reynolds number of 17500.
 Here, the channel is divided into $2000\times 500\times 200$ grid points and has an average spatial resolution of $\Delta x=\Delta y=\Delta z=5.5\eta$.
 The odor source has a Gaussian profile centered at [$825\eta$, $1375\eta$, $550\eta$] with a standard deviation of $3\eta$.
-$\rho(\frac{\partial\bm{u}}{\partial t}+\bm{u}\cdot\nabla\bm{u})$ $=-\nabla P+\mu\nabla^{2}\bm{u}+\bm{f};$ (1)
-$\nabla\cdot\bm{u}$ $=0.$
-$\frac{\partial c}{\partial t}+\mathbf{u}\cdot\nabla c=D\nabla^{2}c+s.$ (2)
+For Environments 5 and 6, noise level is 0.01% relative to concentration at the source.
+See Table for a summary of parameters and how they match the physical dimensions of the domain.
+Table 3: Parameters of the simulations.
+From Left to Right: Simulation ID (1, 2, 3); Length $L$, width $W$, height $H$ of the computational domain; mean horizontal speed $U_{b}=\langle u\rangle$; Kolmogorov length scale $\eta=(\nu^{3}/\epsilon)^{1/4}$ where $\nu$ is the kinematic viscosity and $\epsilon$ is the energy dissipation rate; mean size of gridcell $\Delta x$; Kolmogorov timescale $\tau_{\eta}=\eta^{2}/\nu$; energy dissipation rate $\epsilon=\nu/2\langle(\partial u_{i}/\partial x_{j}+\partial u_{j}/\partial x_{i})^{2}\rangle$; wall unit $y^{+}=\nu/u_{\tau}$ where $u_{\tau}$ is the friction velocity; bulk Reynolds number $Re_{b}=U_{b}(H/2)/\nu$ based on the bulk speed $U_{b}$ and half height; magnitude of velocity fluctuations $u^{\prime}$ relative to the bulk speed; large eddy turnover time $T=H/2u^{\prime}$; $\omega_{\text{save}}$ frequency at which odor snapshots are saved ×$\times\tau_{\eta}$.
+For each simulation, first row reports results in non dimensional units.
+Second and third rows provide an idea of how non-dimensional parameters match dimensional parameters in real flows in air and water, assuming the Kolmogorov length is 1.5 mm in air and 0.4 mm in water.
+Sim ID $L$ $W$ $H$ $U_{b}$ $\eta$ $\Delta x$ $\tau_{\eta}$ $\epsilon$ $y^{+}$ Reb $\frac{u^{\prime}}{U_{b}}$ $T$ $\omega_{\text{save}}\times\tau_{\eta}$
+1 40 8 4 23 0.006 0.025 0.01 39 0.0035 11500 15% $64\tau_{\eta}$ 1
+air 9.50 m 1.90 m 0.96 m 36 $\frac{\text{cm}}{\text{s}}$ 0.15 cm 0.6 cm 0.15 s 6.3e-4 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.09 cm
+water 2.66 m 0.53 m 0.27 m 8.6 $\frac{\text{cm}}{\text{s}}$ 0.04 cm 0.2 cm 0.18 s 3e-5 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.02 cm
+2 20 5 2 14 0.004 0.02 0.005 163 0.0038 7830 15% $95\tau_{\eta}$ 5
+air 7.50 m 1.875 m 0.75 m 17.5 $\frac{\text{cm}}{\text{s}}$ 0.15 cm 0.75 cm 0.15 s 6.3e-4 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.142 cm
+water 2.00 m 0.50 m 0.20 m 3.9 $\frac{\text{cm}}{\text{s}}$ 0.04 cm 0.2 cm 0.18 s 3e-5 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.038 cm
+3 20 5 2 22 0.0018 0.01 0.0025 204 0.0012 17500 13% $141\tau_{\eta}$ 2.5
+air 16.7 m 4.18 m 1.67 m 30.6 $\frac{\text{cm}}{\text{s}}$ 0.15 cm 0.83 cm 0.15 s 6.3e-4 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.1 cm
+water 4.44 m 1.11 m 0.44 m 6.8 $\frac{\text{cm}}{\text{s}}$ 0.04 cm 0.22 cm 0.18 s 3e-5 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.03 cm
+$\rho(\frac{\partial\boldsymbol{u}}{\partial t}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})$ $=-\nabla P+\mu\nabla^{2}\boldsymbol{u}+\boldsymbol{f};$ (1)
+$\nabla\cdot\boldsymbol{u}$ $=0.$
+$\frac{\partial z}{\partial t}+\mathbf{u}\cdot\nabla z=D\nabla^{2}z+s.$ (2)
 Olfactory states, Features & Discretization.
-Each agent stores the odor concentrations detected in the previous $T$ time steps in a vector $\mathbf{M}$. We introduce an adaptive sensitivity threshold function $s_{\text{thr}}(\cdot)$ defined as
+Each agent stores the odor concentrations detected in the previous $T$ time steps in a vector $\mathbf{M}=(z(s(t-T),t-T),...,z(s(t),t))$.
+We introduce an adaptive sensitivity threshold function $s_{\text{thr}}(\cdot)$ defined as
 $s_{\text{thr}}(T):=\max\Big{\{}\frac{C_{\text{thr}}}{T}\sum\limits_{i=1}^{T}M_{i},n_{\text{thr}}\Big{\}},$ (3)
@@
 where $e_{i}$ denotes the $i$-th canonical base.
+Actions are steps in any of the 4 directions, labeled relative to the mean flow which assumed fixed and known.
+The gridworld is infinite, in that agents can leave indefinitely.
+If they exit the simulation box, they continue to receive zero signal and negative reward -0.001.
 As explained in the main text, actions are selected using one of two policies according to the current olfactory state $o_{t}$.
@@
 Note that this strategy only provides exploration after the backtracking fails to recover detections. Also, if agents start with no detection at time 0, the procedure is equivalent to Brownian motion.
+Circling recovery.
+In this case, the recovery strategy consists of adopting a circling behavior Stupski and van Breugel (2024) when the agent is in a void state.
+The agent keeps in memory two counters, $t_{\text{void}}$ and $T_{\text{change}}$, as well as an action $a_{\text{void}}\in\mathcal{A}$, initialized to 00, $1$, and $e_{1}$, respectively.
+The first counter represents the consecutive number of void observations, $a_{\text{void}}$ is the action to play when the void state is reached, and $T_{\text{change}}$ is a time threshold that indicates when to switch the action $a_{\text{void}}$.
+When the agent reaches a void state, it plays the action $a_{\text{void}}$ and increments the counter $t_{\text{void}}$ by one.
+If $t_{\text{void}}=T_{\text{change}}$, then the agent resets $t_{\text{void}}$ to zero, increases $T_{\text{change}}$ by one, and updates the action $a_{\text{void}}$ as follows
+$a_{\text{void}}\leftarrow\left\{\begin{array}[]{cc}e_{2}&\text{if }a_{\text{void}}=e_{1}\\ -e_{1}&\text{if }a_{\text{void}}=e_{2}\\ -e_{2}&\text{if }a_{\text{void}}=-e_{1}\\ e_{1}&\text{if }a_{\text{void}}=-e_{2}\\ \end{array}\right.$
+When the agent receives a non-void observation, it resets $t_{\text{void}}$ to 00, $T_{\text{change}}$ to $1$, and $a_{\text{void}}$ to $e_{1}$.
+Cast-Surge recovery.
+In this case, the agent plays a cast-surge behavior Baker (1990) when reaching the void state.
+As in the circling recovery strategy, the agent keeps two counters, $t_{\text{void}}$ and $T_{\text{change}}$, as well as an action $a_{\text{void}}\in\mathcal{A}$, initialized to 00, $1$, and $e_{2}$, respectively.
+At every step in the void state, the agent plays the action $a_{\text{void}}$ and increments the void counter $t_{\text{void}}$ by one.
+If $t_{\text{void}}=T_{\text{change}}$, the agent takes an upwind step, doubles the value of $T_{\text{change}}$, updates $a_{\text{void}}$ by setting it to $-a_{\text{void}}$, and resets $t_{\text{void}}$ to zero.
+When the agent receives a non-void observation, it resets $t_{\text{void}}$ to 00, $T_{\text{change}}$ to $1$, and $a_{\text{void}}$ to $e_{2}$.
 Learned recovery.
@@
 For each admissible location $(i,j)$, we define four performance metrics:
-$G(i,j)=$ $\langle e^{-\lambda\tau(i,j)}-\frac{\eta}{1-\gamma}(1-e^{-\lambda\tau(i,j)})\rangle$
+$G(i,j)=$ $\langle e^{-\lambda\tau(i,j)}-\frac{\sigma}{1-\gamma}(1-e^{-\lambda\tau(i,j)})\rangle$
 $f^{+}(i,j)=$ $\frac{n_{\text{success}}(i,j)}{n_{\text{reps}}}$
@@
 […]
+4.1 Acknowledgements
+This research was supported by grants to AS from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101002724 RIDING), the Air Force Office of Scientific Research under award number FA8655-20-1-7028, and the National Institutes of Health (NIH) under award number R01DC018789.
+L.
+R.
+and M.R.
+acknowledge the financial support of the European Research Council (grant SLING 819789), the European Commission (Horizon Europe grant ELIAS 101120237), the US Air Force Office of Scientific Research (FA8655-22-1-7034), the Ministry of Education, University and Research (FARE grant ML4IP R205T7J2KP; grant BAC FAIR PE00000013 funded by the EU - NGEU) and the Center for Brains, Minds and Machines (CBMM), funded by NSF STC award CCF-1231216.
+M.R.
+is a member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM).
+This work represents only the view of the authors.
+The European Commission and the other organizations are not responsible for any use that may be made of the information it contains.
+We thank Francesco Viola for support and discussions regarding computational fluid dynamics as well as Antonio Celani, Venkatesh Murthy, Yujia Qi, Francesco Boccardo, Luca Gagliardi, Francesco Marcolli and Arnaud Ruymaekers for comments on the manuscript.
+Data Availability.
+The datasets and code used to perform the experiments are available at the following links